### v1

# DIRECT: Deep Active Learning under Imbalance and Label Noise

Abstract
Class imbalance is a prevalent issue in real world machine learning applications, often leading to poor performance in rare and minority classes. With an abundance of wild unlabeled data, active learning is perhaps the most effective technique in solving the problem at its root – collecting a more balanced and informative set of labeled examples during annotation. In this work, we propose a novel algorithm that first identifies the class separation threshold and then annotate the most uncertain examples from the minority classes, close to the separation threshold. Through a novel reduction to one-dimensional active learning, our algorithm DIRECT is able to leverage the classic active learning literature to address issues such as batch labeling and tolerance towards label noise. Compared to existing algorithms, our algorithm saves more than 15% of the annotation budget compared to state-of-art active learning algorithm and more than 90% of annotation budget compared to random sampling.
Machine Learning, ICML
1 Introduction
Large-scale deep learning models is playing an increasing role across many industries. Human feedback and annotations have played a significant role in developing such systems. Increasingly, we believe the role of human would shift to annotating rare yet important cases such as safety risks. In this paper, we propose an active learning algorithm to address the class imbalance problem, where we sequentially and adaptively choose examples for annotation. Our proposed algorithm optimizes for the class-balancedness and informativeness of the annotation collected.
Inspired by previous work by Zhang et al. (2022), we propose a novel reduction of the imbalanced classification problem into an one-dimensional active learning problem by arranging unlabeled examples into a list ordered by uncertainty scores. With this reduction, we apply classic active learning literature in finding the optimal separation threshold between the minority and majority classes. We then annotate examples near this threshold to gather uncertain examples.
On the other hand, existing state-of-art algorithm like GALAXY (Zhang et al., 2022) does not allow parallel annotation, and performs poorly when label noise is high. With our novel one-dimensional reduction, we inherit the noise tolerance and batch labeling advantages of classic active learning algorithms, allowing our deep active learning algorithm DIRECT to overcome such issues.
To summarize our main contributions:
  • We propose a novel reduction that bridges the advancement in theoretical active learning literature to imbalanced active classification.
  • Our novel algorithm addresses the prevalent imbalancedness issue by annotating a more class-balanced and informative set.
  • Compared to state-of-art algorithm GALAXY (Zhang et al., 2022), our algorithm allows parallel annotation and is tolerant to label noise. It also saves $15\%$ more annotation cost compared to GALAXY, and $90\%$ of the annotation cost compared to random sampling.
2 Related Work
Deep Active Learning Under Class Balance Active Learning strategies sequentially and adaptively choose examples for annotation (Settles, 2009). Many uncertainty methods active learning methods extend the traditional active learning literature such as margin, least confidence and entropy sampling (Tong & Koller, 2001; Settles, 2009; Balcan et al., 2006; Kremer et al., 2014). These methods have also been shown to perform among top methods when fine-tuning large pretrained models and combined with semi-supervised learning algorithms (Zhang et al., 2023a). More sophisticated methods have also been proposed to optimize the chosen unlabeled examples’ uncertainty (Gal et al., 2017; Ducoffe & Precioso, 2018; Beluch et al., 2018), diversity (Sener & Savarese, 2017; Geifman & El-Yaniv, 2017; Citovsky et al., 2021), or a mix of both (Ash et al., 2019, 2021; Wang et al., 2021; Elenter et al., 2022; Mohamadi et al., 2022). However, these methods often perform poorly under prevalent and realistic scenarios such as under label noises (Khosla et al., 2022) or under class imbalance (Kothawade et al., 2021; Zhang et al., 2022, 2023a).
Deep Active Learning under Imbalance Data imbalance and rare instances are prevalent in almost all modern machine learning applications. Active learning techniques are effective in addressing the problem in its root by collecting a more class-balanced dataset (Aggarwal et al., 2020; Kothawade et al., 2021; Emam et al., 2021; Zhang et al., 2022; Coleman et al., 2022; Jin et al., 2022; Cai, 2022; Zhang et al., 2023b). To this end, Kothawade et al. (2021) propose a submodular-based method that actively annotates examples similar to known examples of rare instances. Our work is motivated by Zhang et al. (2022), where they propose GALAXY, an algorithm that construct one-dimensional linear graphs and apply graph-based active learning techniques in annotating a set of examples that are both class-balanced and uncertain. While GALAXY outperforms existing algorithms, due to a bisection procedure involved, it does not allow parallel annotation. In addition, bisection procedures are generally not robust against label noises, a prevalent challenge in real world annotation tasks. Our algorithm DIRECT mitigates all of the above shortcomings of GALAXY while outperforming it even with synchronous labeling and no label noise, beating GALAXY in its own game.
Lastly, we distinguish our work from Zhang et al. (2023b), where the paper studies the algorithm selection problem. Unlike our goal of proposing a new deep active learning algorithm, the paper proposes meta algorithms to choose the right active learning algorithm among a large number of candidate algorithms.
3 Preliminary
3.1 Notations
We study the pool-based active learning problem, where an initial unlabeled set of $N$ examples $X=\{x_{1},...,x_{N}\}$ are available for annotation. We let $\mathcal{X}$ denote the space of possible examples. Their corresponding labels $Y=\{y_{1},...,y_{N}\}$ are initially unknown. Furthermore, we study the multi-class classification problem, so the space of labels $\mathcal{Y}:=[K]$, where each element denotes the class index and $K$ is the total number of classes. Moreover, let $N_{1},...,N_{K}$ denote the number of examples in $X$ in each class. We define the imbalance ratio as $\gamma=\frac{\min_{k\in[K]}N_{k}}{\max_{k\in[K]}N_{k}}$.
A deep active learning algorithm iteratively chooses batches of examples for annotation. During the $t$-th iteration, the algorithm is given labeled and unlabeled sets of examples, $L_{t}$ and $U_{t}$ respectively, where $L_{t}\cup U_{t}=X$ and $L_{t}\cap U_{t}=\emptyset$. The algorithm then chooses $B$ examples from the unlabeled set $X^{(t)}\subseteq U_{t}$ for their corresponding labels $Y^{(t)}$ to be annotated. The labeled and unlabeled sets are then updated, i.e., $L_{t+1}\leftarrow L_{t}\cup X^{(t)}$ and $U_{t+1}\leftarrow U_{t}\backslash X^{(t)}$. Based on new labeled set $L_{t+1}$ and its corresponding labels, a neural network $f_{t}:\mathcal{X}\rightarrow\mathcal{Y}$ is trained to inform the choice for the next iteration.
The ultimate goal of deep active learning is to obtain high predictive accuracy for the trained neural network while annotating as few examples as possible.
3.2 Review of Imbalanced Active Learning: GALAXY
We first consider a binary case of imbalanced classification, where we assume $N_{1}<N_{2}$ without loss of generality. After a total of $T$ rounds and sampling $TB$ examples, random sampling would annotate a subset of $X$ with an imbalance ratio close to $\frac{N_{1}}{N_{2}}$. On the other hand, as shown previously in Zhang et al. (2022), uncertainty based methods such as confidence (Settles, 2009), margin (Tong & Koller, 2001; Balcan et al., 2006) and entropy (Kremer et al., 2014) sampling could annotate a set of examples with imbalance ratio arbitrarily low. The reasoning is simply demonstrated in Figure 0(a), where choosing examples with predictive sigmoid score $\widehat{p}$ closest to .5.5.5.5 may end up annotating examples (all) from the majority class. It is also empirically shown that annotating examples most certain to be in the minority class is not the most effective in improving model performance.
[…]
(b) GALAXY spends approximately equal annotation budget around both cuts, while the cut on the right would yield examples mostly in the majority class.
Figure 1: Figures demonstrating GALAXY. Ordered lists of examples are ranked by the predictive sigmoid score $\widehat{p}$. The ground truth label of each example is represented by its border – solid blue lines for class $1$ and dotted red lines for class $2$. Annotated examples are shaded.
Therefore, the key objective in imbalanced active learning is labeling examples that are both uncertain and class balanced. The proposed algorithm GALAXY (Zhang et al., 2022) follows the following two-phased procedure: 1. Identify the optimal separation threshold $\widehat{p}^{\star}$, as shown in Figure 0(a). 2. Annotate examples in the minority class closest to $\widehat{p}^{\star}$.
GALAXY draws inspiration from graph-based active learning in identifying $\widehat{p}^{\star}$. In particular, it follows a modified bisection procedure that sequentially annotate examples to search for all cuts, namely thresholds with a class 1 example to the left and a class 2 example to the right (see Figure 0(b)). GALAXY then annotate examples close to all such cuts. However, the algorithm suffers from two weaknesses. Firstly, a bisection procedure only allows sequential labeling, which prevents multiple annotators labeling in parallel. Secondly, the problem of identifying the optimal separation threshold is a noisy procedure, since the predictive sigmoid scores may not perfectly separate the minority class examples from the majority. GALAXY mitigates the problem by wastefully identifying and annotating around all cuts, which could lead to annotating large number of majority class examples (shown in Figure 0(b)).
4 A Robust Algorithm with Batch Labeling
In this section, we formally define optimal separation threshold and pose the problem of identifying it as an 1-dimensional reduction to the agnostic active learning problem. We then propose a novel algorithm inspired by the agnostic active learning literature (Balcan et al., 2006; Dasgupta et al., 2007; Hanneke et al., 2014; Katz-Samuels et al., 2021).
4.1 An 1-D Reduction to Agnostic Active Learning
We start by considering the imbalanced binary classification setting mentioned in Section 3.2. When given a neural network model, we let $\widehat{p}:\mathcal{X}\rightarrow[0,1]$ be the predictive function mapping examples to sigmoid scores. Here, a higher sigmoid score represents a higher confidence of the example being in class $2$. Let $q_{1},...,q_{N}$ denote the sigmoid scores of every example in the pool $X$, where $q_{i}=\widehat{p}(x_{i})$ for each $i\in[N]$. We also sort examples by the sigmoid predictive score similar to Section 3.2. Formally, we let $q_{(1)}\leq q_{(2)}\leq\cdots q_{(N)}$ be a sorted permutation of $q_{1},...,q_{N}$, and let $y_{(1)},y_{(2)},...,y_{(N)}$ denote the sorted list’s corresponding labels. We are now ready to define the optimal separation threshold as described in Section 3.2 below.
Definition 4.1.
Given the sorted predictive sigmoid scores $0\leq q_{(1)}\leq\cdots\leq q_{(N)}\leq 1$, we define the optimal separation threshold as $j^{\star}\in[N]$ such that
$j^{\star}=\operatorname*{arg\,max}_{j}$ (|$\left(\lvert\{y_{(i)}=1:1\leq i\leq j\}\rvert\right.$
|$\left.-\lvert\{y_{(i)}=2:1\leq i\leq j\}\rvert\right).$ (1)
In other words, $j^{\star}$ has the largest difference in the number of class $1$ examples from class $2$ examples to its left. Furthermore, ties are broken by choosing the largest $j^{\star}$ that attains the argmax if class $1$ is the minority class and the lowest $j^{\star}$ otherwise.
1D Reduction. We now provide a reduction of finding $j^{\star}$ to an 1-dimensional agnostic active learning problem. In particular, let =$\mathcal{H}=\{h_{0},h_{1},...,h_{N}\}$ be the hypothesis class. Here each hypothesis $h_{j}:[0,1]\rightarrow\{1,2\}$ maps sigmoid scores to binary target classes and
$h_{j}(q)=\begin{cases}1&\text{if}\,q\leq q_{(j)}\\ 2&\text{if}\,q>q_{(j)}\end{cases}.$
The empirical zero-one loss for each hypothesis is then defined as $\mathcal{L}(h_{j})=\sum_{i=1}^{N}\mathbf{1}\{h_{j}(q_{(i)})\neq y_{(i)}\}$. In Appendix Appendix A., we show that optimizing for the zero-one loss $\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$ is equivalent to (1). Namely, $j^{\star}\in\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$, transforming the problem of finding the optimal separation threshold to an 1-D agnostic active learning problem.
Multi-Class Classification. To generalize the above problem formulation to multi-class classification, we follow a similar strategy to Zhang et al. (2022). Specifically, for each class $k$, one can view the problem of class-$k$ v.s. others as a binary classification problem. In other words, class $k$ can be seen as class $1$ and classes $\{1,...,k-1,k+1,...,N\}$ can be seen as class $2$. The goal therefore becomes finding all $K$ optimal separation thresholds, which is equivalent with solving $K$ 1-D agnostic active learning problems. Moreover, let $\hat{p}:\mathcal{X}\rightarrow\Delta^{(K-1)}$ denote the neural network prediction function mapping examples to sigmoid scores. Instead of using sigmoid scores, for class $k$, we use the margin scores $q_{i}^{k}:=\max_{k^{\prime}}[\hat{p}(x_{i})]_{k^{\prime}}-[\hat{p}(x_{i})]_{k}$ to sort the examples. We also note that sorting with margin scores is equivalent to sorting with sigmoid scores in the binary classification setting, so this is a strictly more general formulation to multi-class classification.
4.2 Algorithm
We are now ready to state our algorithm as shown in Algorithm 1. The algorithm follows a two-phased, where the first phase aims to identify the optimal separation threshold for each class. The second phase then annotates examples closest to the estimated optimal separation thresholds for each class.
  Input: Pool $X$, number of rounds $T$, retraining batch size $B_{\text{train}}$, number of parallel annotations $B_{\text{parallel}}$.
  Initialize: Uniformly sample $B$ elements without replacement from $X$ to form $L_{0}$. Let $U_{0}\leftarrow X\backslash L_{0}$.
  for $t=1,...,T-1$ do
     Train neural network on $L_{t-1}$ and obtain $f_{t-1}$.
     Find optimal separation thresholds
     Initialize labeled set $L_{t}\leftarrow L_{t-1}$.
     Initialize budget per class $b\leftarrow\frac{B_{\text{train}}}{2K}$.
     for $k$ in $\text{RandPerm}(\{1,...,K\})$ do
        Compute margin scores $q_{1}^{k},...,q_{N}^{k}$ based on $f_{t-1}$ and sort them by $q_{(1)}^{k}\leq\cdots\leq q_{(N)}^{k}$.
        Find optimal separation threshold: $L_{t}\leftarrow\text{VReduce}(L_{t},b,k,B_{\text{parallel}},\{(x_{(i)},y_{(i)})\}_{i=1}^{N})$.
     end for
     Annotate rare and uncertain examples
     Compute budget per class $b\leftarrow\frac{B_{\text{train}}-|L_{t}|}{K}$.
     for $k$ in $\text{RandPerm}(\{1,...,K\})$ do
        Estimate separation threshold
$\widehat{j}^{k}\leftarrow$ $\operatorname*{arg\,max}_{j}(|\{y_{(i)}=1:x_{(i)}\in L_{t}\,\text{and}\,i\leq j\}|$
|$-|\{y_{(i)}=1:x_{(i)}\in L_{t}\,\text{and}\,i\leq j\}|)$
and break ties by choosing the index closest to $\frac{N}{2}$. Annotate $b$ unlabeled examples with sorted indices closest to $\widehat{j}^{k}$ and insert to $L_{t}$.
     end for
  end for
  Return: Train final classifier $f_{T}$ based on $L_{T}$.
Algorithm 1 DIRECT: DImension REduction for aCTive Learning under Imbalance and Label Noise
To identify the optimal separation threshold, we loop over each class $k$ and run a fix-budget batch active learning procedure for the one-dimensional reduction (as shown in Algorithm 2). Specifically, we keep a version space of the hypotheses with higher likelihood of being the optimal separation threshold. We annotate examples sequentially in batches and shrink the version space by a fixed rate each iteration. The shrinkage rate $c$ is determined by the budget and batch size, so that after the final iteration, the version space has exactly one hypothesis left.
To address batch labeling, we note the number of examples one collect before retraining is usually far greater than the number of annotators annotating in parallel. In particular, we let $B_{\text{train}}$ denote the number of examples the algorithm collects before the neural network is retrained. In practice, this number is usually determined by the practical constraints of computational training cost. On the other hand, we let $B_{\text{parallel}}$ denote the number of examples annotated in parallel. As discussed above, $B_{\text{parallel}}<<B_{\text{train}}$ in practice. Lastly, as will be discussed in Section 6, a variant of our algorithm can also be proposed for asynchronous labeling.
  Input: Labeled set $L$, budget $b$, class of interest $k$, parallel batch size $B_{\text{parallel}}$, examples and ground truth labels sorted by uncertainty $\{(x_{(i)},y_{(i)})\}_{i=1}^{N}$ (ground truth labels are hidden to the learner).
  Initialize: Version space: shortest segment of indices $[I,J]$, where $X_{I},X_{J}\in L$, $\forall i\leq I,x_{(i)}\in L\implies y_{(i)}=k$ and $\forall j\geq J,x_{(j)}\in L\implies y_{(j)}\neq k$.
  Initialize: Number of iterations $m\leftarrow\frac{b}{B_{\text{parallel}}}$. Shrinking factor $c\leftarrow\sqrt[m]{J-I}$
  for $t=1,...,m$ do
     Sample uniformly at random $B_{\text{parallel}}$ unlabeled examples in $x_{(I)}$, …, $x_{(J)}$ for annotation and insert to $L$.
     Find $[I^{\prime},J^{\prime}]\subset[I,J]$ such that
$I^{\prime},J^{\prime}=\operatorname*{arg\,min}_{i,j:j-i=\frac{1}{c}(J-I)}\max\{\widehat{\mathcal{L}}^{k}(i),\widehat{\mathcal{L}}^{k}(j)\}\quad\text{where }$
^$\widehat{\mathcal{L}}^{k}(s)=\sum_{\begin{subarray}{c}r\leq s:\\ x_{(r)}\in L\end{subarray}}\mathbf{1}\{y_{(r)}\neq k\}+\sum_{\begin{subarray}{c}r>s:\\ x_{(r)}\in L\end{subarray}}\mathbf{1}\{y_{(r)}=k\}.$
     Update version space $I,J\leftarrow I^{\prime},J^{\prime}$.
  end forReturn: Updated labeled set $L$.
Algorithm 2 VReduce: Version Space Reduction
5 Experiments
(a) Imbalanced CIFAR-10, two classes
(b) Imbalanced CIFAR-10, three classes
(c) Imbalanced CIFAR-100, two classes
(d) Imbalanced CIFAR-100, three classes
(e) Imbalanced SVHN, two classes
(f) Imbalanced SVHN, three classes
Figure 2: Performance of DIRECT vs different baselines across six different dataset settings.
5.1 Experiment Setup
Our experiment setup follows closely from Zhang et al. (2022). We use the CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009) and SVHN (Netzer et al., 2011) image classification datasets. The original CIFAR-10 and SVHN datasets have 10 classes while the CIFAR-100 dataset has 100 classes. We construct an extremely imbalanced dataset by grouping a large number of classes into one majority class. For example, suppose we have a dataset with $M$ classes. We generate an imbalanced dataset with $K$ classes ($K<M$) by reusing classes $11,...K-$ 1 from the original dataset and combining the rest of the classes $K,...,M$ into a single majority class $K$. Detailed imbalance ratios are shown in Table 1.
Name $K$ $N_{K}$ $\sum_{k=1}^{K-1}N_{k}$ Imb Ratio $\gamma$
CIFAR-10 $2$ $45000$ $5000$$1$.1111.1111.1111
CIFAR-10 $3$ $40000$ $10000$1250.1250.1250.1250
CIFAR-100 $2$ $49500$ $500$0101.0101.0101.0101
CIFAR-100 $3$ $49000$ $1000$0102.0102.0102.0102
SVHN $2$ $68309$ $4948$0724.0724.0724.0724
SVHN $3$ $54448$ $18809$2546.2546.2546.2546
Table 1: Dataset setting for extremely unbalanced scenario. $N_{K}$ denotes the number of images in the majority class while $\sum_{k=1}^{K-1}N_{k}$ is the total number of images in all minority classes. $\epsilon$ is the class imbalance factor defined in Section 3.1.
We compare our algorithm DIRECT against nine baselines: GALAXY (Zhang et al., 2022), SIMILAR (Kothawade et al., 2021), BADGE (Ash et al., 2019), BASE (Emam et al., 2021), BAIT (Ash et al., 2021), Cluster Margin (Citovsky et al., 2021), Confidence Sampling (Settles, 2009), Most Likely Positive (Jiang et al., 2018; Warmuth et al., 2001, 2003) and Random Sampling. The highlight surrounding the baselines represents the standard error plotted as the confidence intervals. Due to computational constraints, we only have single runs for SIMILAR, BADGE, and BAIT. For all other baselines, we average over four runs.
For all experiments, we use PyTorch’s ResNet-18 model pretrained on ImageNet for initialization. We use the Adam optimization algorithm with 500 epochs for each L and the learning rate set to 0.01.
5.2 Results
Synchronous Labeling ($B_{\text{parallel}}=1$). In our first set of experiments, we compare DIRECT against GALAXY in the synchronous labeling setting by setting $B_{\text{parallel}}=1$ in our algorithm. As shown in Figure 2(a), DIRECT outperforms GALAXY by saving an additional $15\%$ of the labeling budget. Compared to random sampling, DIRECT is able to around $90\%$ of annotation cost.
Batch Labeling. We observe that on all six datasets shown in Figure 2, DIRECT performs comparably to GALAXY even when under parallel annotation ($B_{\text{parallel}}=5$). However, we do observe a slight drop in label-efficiency when comparing to synchronous labeling.
6 Future Work
In this paper, we addressed the batch sampling problem of GALAXY by annotating multiple examples in parallel. However, in practice, examples are often annotated asynchronously. We believe this is a valuable future direction and a potential solution is to utilize an asynchronous variant of one-dimensional active learning algorithm.
Acknowledgements
This work has been supported in part by NSF Award 2112471.
Appendix A.
Lemma .1.
The agnostic active learning reduction is equivalently finding the optimal separation threshold. Namely,
$\operatorname*{arg\,min}_{j}\mathcal{L}(h_{j})=\operatorname*{arg\,max}_{j}\left(\lvert\{y_{(i)}=1:1\leq i\leq j\}\rvert-\lvert\{y_{(i)}=2:1\leq i\leq j\}\rvert\right)$
Proof.
Recall the definitions: $h_{j}(q)=\begin{cases}1&\text{if}\,q\leq q_{(j)}\\ 2&\text{if}\,q>q_{(j)}\end{cases}$ and $\mathcal{L}(h_{j})=\sum_{i=1}^{N}\mathbf{1}\{h_{j}(q_{(i)})\neq y_{(i)}\}$, we can expand the loss as follows
$\operatorname*{arg\,min}_{j}\mathcal{L}(h_{j})$ $=\operatorname*{arg\,min}_{j}\sum_{i=1}^{N}\mathbf{1}\{h_{j}(q_{(i)})\neq y_{(i)}\}$
$=\operatorname*{arg\,min}_{j}N-\sum_{i=1}^{N}\mathbf{1}\{h_{j}(q_{(i)})=y_{(i)}\}$
$=\operatorname*{arg\,max}_{j}\sum_{i=1}^{N}\mathbf{1}\{h_{j}(q_{(i)})=y_{(i)}\}$
$=\operatorname*{arg\,max}_{j}\left(\sum_{i=1}^{j}\mathbf{1}\{y_{(i)}=1\}\right)+\left(\sum_{i=j+1}^{N}\mathbf{1}\{y_{(i)}=2\}\right)$
$=\operatorname*{arg\,max}_{j}\left(\sum_{i=1}^{j}\mathbf{1}\{y_{(i)}=1\}\right)+\left(\sum_{i=j+1}^{N}\mathbf{1}\{y_{(i)}=2\}\right)-\left(\sum_{i=1}^{N}\mathbf{1}\{y_{(i)}=2\}\right)$
$=\operatorname*{arg\,max}_{j}\sum_{i=1}^{j}\left(\mathbf{1}\{y_{(i)}=1\}-\mathbf{1}\{y_{(i)}=2\}\right)$
lt="[LOGO]">

### v2

@@
 With an abundance of wild unlabeled data, active learning is perhaps the most effective technique in solving the problem at its root – collecting a more balanced and informative set of labeled examples during annotation.
-In this work, we propose a novel algorithm that first identifies the class separation threshold and then annotate the most uncertain examples from the minority classes, close to the separation threshold.
+Label noise is another common issue in data annotation jobs, which is especially challenging for active learning methods.
+In this work, we conduct the first study of active learning under both class imbalance and label noise.
+We propose a novel algorithm that robustly identifies the class separation threshold and annotates the most uncertain examples that are closest from it.
 Through a novel reduction to one-dimensional active learning, our algorithm DIRECT is able to leverage the classic active learning literature to address issues such as batch labeling and tolerance towards label noise.
-Compared to existing algorithms, our algorithm saves more than 15% of the annotation budget compared to state-of-art active learning algorithm and more than 90% of annotation budget compared to random sampling.
+We present extensive experiments on imbalanced datasets with and without label noise.
+Our results demonstrate that DIRECT can save more than 60% of the annotation budget compared to state-of-art active learning algorithms and more than 80% of annotation budget compared to random sampling.
 Machine Learning, ICML
 1 Introduction
-Large-scale deep learning models is playing an increasing role across many industries.
-Human feedback and annotations have played a significant role in developing such systems.
-Increasingly, we believe the role of human would shift to annotating rare yet important cases such as safety risks.
-In this paper, we propose an active learning algorithm to address the class imbalance problem, where we sequentially and adaptively choose examples for annotation.
-Our proposed algorithm optimizes for the class-balancedness and informativeness of the annotation collected.
+Large-scale deep learning models are playing increasingly important roles across many industries.
+Human feedback and annotations have played a significant role in developing such systems through.
+Progressively over time, we believe the role of humans in a machine learning pipeline would shift to annotating rare yet important cases.
+However, under data imbalance, the typical strategy of randomly choosing examples for annotation becomes especially inefficient.
+This is because the majority of the labeling budget would be spent on common and well-learned classes, resulting in insufficient rare class examples for training an effective model.
+For many annotation jobs, this challenge of data imbalance is further compounded by label noises induced from annotators.
+In this paper, we mitigate these issues by proposing a deep active learning strategy.
+Our algorithm DIRECT sequentially and adaptively chooses informative and more class-balanced examples for annotation while being robust to noisy annotations.
+To the best of our knowledge, this is the first deep active learning study to address the challenging yet prevalent scenario where both imbalance and label noise coexist.
 Inspired by previous work by Zhang et al.
-(2022), we propose a novel reduction of the imbalanced classification problem into an one-dimensional active learning problem by arranging unlabeled examples into a list ordered by uncertainty scores.
-With this reduction, we apply classic active learning literature in finding the optimal separation threshold between the minority and majority classes.
-We then annotate examples near this threshold to gather uncertain examples.
-On the other hand, existing state-of-art algorithm like GALAXY (Zhang et al., 2022) does not allow parallel annotation, and performs poorly when label noise is high.
-With our novel one-dimensional reduction, we inherit the noise tolerance and batch labeling advantages of classic active learning algorithms, allowing our deep active learning algorithm DIRECT to overcome such issues.
+(2022), we propose a novel reduction of the imbalanced classification problem into one-dimensional active learning problems.
+For each class, our reduction sorts unlabeled examples into an list ordered by one-vs-rest margin scores.
+The goal of DIRECT is to find the optimal separation threshold which best separates the examples in any given class from the rests.
+With an equivalence to finding the best threshold classifier, we were able to employ classic active learning literature for obtaining the separation threshold.
+Once the optimal separation threshold is identified, DIRECT annotates a set of more class-balanced examples near this threshold.
+As examples around the optimal separation threshold tend to have high uncertainty, DIRECT annotates examples that are both class-balanced and informative.
+With the reduction to classic one-dimensional active learning literature, DIRECT also inherits the advantages from classic active learning algorithms such as robustness against label noise and batch labeling.
+This is in contrast to the state-of-art imbalanced active learning algorithm GALAXY proposed by Zhang et al.
+(2022), which requires sequential labeling and is much less robust against label noise.
+As demonstrated in Figure 1, under both noisy and noiseless label settings, DIRECT improves significantly over GALAXY’s label efficiency – DIRECT requires less annotations to achieve the same accuracy.
+It achieves state-of-art label efficiency gain on both supervised fine-tuning of ResNet-18 and semi-supervised fine-tuning of large pretrained model under the LabelBench framework (Zhang et al., 2023a).
+Lastly, DIRECT also allows for parallel annotations with various batch sizes.
+(a) Imbalanced CIFAR-10, two classes, no label noise.
+(b) Imbalanced CIFAR-100, two classes, 20% label noise.
+(c) LabelBench FMoW (62 classes with imbalance), no label noise
+Figure 1: Performance of DIRECT over existing baselines for both noiseless and noisy settings.
+The x-axis represents the total number of labeled examples so far and the y-axis shows the neural network’s balanced accuracy.
+Both (a) and (b) are using supervised training of ResNet-18.
+In (c), we finetune CLIP ViT-B32 model in combination of semi-supervised training under the LabelBench framework (Zhang et al., 2023a).
+$B_{\text{parallel}}$ is the batch size indicating the number of parallel annotators for the annotation job.
+$B_{\text{parallel}}=1$ indicates the synchronous annotation requirement by GALAXY.
+Our algorithm DIRECT takes pre-specified $B_{\text{parallel}}$ as input, which is determined by real world decisions.
 To summarize our main contributions:
-  • We propose a novel reduction that bridges the advancement in theoretical active learning literature to imbalanced active classification.
-  • Our novel algorithm addresses the prevalent imbalancedness issue by annotating a more class-balanced and informative set.
-  • Compared to state-of-art algorithm GALAXY (Zhang et al., 2022), our algorithm allows parallel annotation and is tolerant to label noise.
-It also saves $15\%$ more annotation cost compared to GALAXY, and $90\%$ of the annotation cost compared to random sampling.
+  • We propose a novel reduction that bridges the advancement in theoretical active learning literature to imbalanced active classification for deep neural networks.
+  • Our algorithm DIRECT addresses the prevalent imbalance and label noise issues and annotates a more class-balanced and informative set of examples.
+  • Compared to state-of-art algorithm GALAXY (Zhang et al., 2022), DIRECT allows parallel annotation by multiple annotators while still maintaining significant label-efficiency improvement.
+  • We conduct experiments across eight dataset settings, four levels of label noise and for both ResNet-18 and large pretrained model (CLIP ViT-B32).
+DIRECT consistently performs the best compared against existing baseline algorithms.
+Quantitatively, DIRECT can save more than $60\%$ annotation cost compared to GALAXY, and more than $80\%$ annotation cost compared to random sampling.
 2 Related Work
-Deep Active Learning Under Class Balance Active Learning strategies sequentially and adaptively choose examples for annotation (Settles, 2009).
-Many uncertainty methods active learning methods extend the traditional active learning literature such as margin, least confidence and entropy sampling (Tong & Koller, 2001; Settles, 2009; Balcan et al., 2006; Kremer et al., 2014).
+Class-Balanced Deep Active Learning Active Learning strategies sequentially and adaptively choose examples for annotation (Settles, 2009).
+Many uncertainty-based deep active learning methods extend the traditional active learning literature such as margin, least confidence and entropy sampling (Tong & Koller, 2001; Settles, 2009; Balcan et al., 2006; Kremer et al., 2014).
 These methods have also been shown to perform among top methods when fine-tuning large pretrained models and combined with semi-supervised learning algorithms (Zhang et al., 2023a).
@@
 Unlike our goal of proposing a new deep active learning algorithm, the paper proposes meta algorithms to choose the right active learning algorithm among a large number of candidate algorithms.
+Deep Active Learning under Label Noise Label noisy settings has rarely been studied in the deep active learning literature.
+Related but tangential to our work, several papers have studied to use active learning for cleaning existing noisy labels (Lin et al., 2016; Younesian et al., 2021).
+In this line of work, they assume access to an oracle annotator that will provide clean labels when queried upon.
+This is fundamentally different from our work, where our annotator may provide noisy labels.
+Another line of more theoretical active learning research studies active learning with multiple annotators with different qualities (Zhang & Chaudhuri, 2015; Chen et al., 2022).
+The primary goal in these work is to identify examples a weak annotator and a strong annotator may disagree, in order to only use the strong annotator on those instances.
+In our work, we assume access to a single source of annotator that is noisy, which is prevalent in annotation jobs today.
+Recently, Khosla et al.
+(2022) proposed a novel deep active learning algorithm specialized for Heteroskedastic noise, where different “regions” of examples are subject to different levels of noise.
+Unlike their work, our work is agnostic to the noise distributions and conduct experiments on uniformly random corrupted labels.
+To the best of our knowledge, no existing deep active learning work has studied the scenario where both imbalance and label noise present.
+Yet, this setting is the most prevalent in real-world annotation applications.
+Agnostic Active Learning A critical part of our solution is the reduction to a one dimensional agnostic active learning problem, e.g.
+Section 4.1.
+There is extensive work on agnostic learning in general, and we refer the interested reader to the survey (Hanneke et al., 2014) for a thorough discussion.
+All of these works, beginning with the seminal works (Balcan et al., 2006; Dasgupta et al., 2007) follow, a familiar paradigm of disagreement based learning.
+This involves maintainin a version space of promising hypothesis at each time, and constructing the disagreement region as a set of unlabeled examples for which there exists two hypotheses that disagree on their predictions.
+The example to label is chosen by sampling from a informative sampling distribution computed over the disagreement region.
+Several approaches have been proposed for computing the sampling distribution, e.g.
+Jain & Jamieson (2019); Katz-Samuels et al.
+(2020, 2021); Huang et al.
+(2015).
+As described in Section 4.2, our main subroutine VReduce is equivalent to fixed-budget one dimensional threshold disagreement learning based on the ACED algorithm of Katz-Samuels et al.
+(2021).
+We remark that the problem of class imbalance has been largely ignored by the classical work on traditional agnostic learning, and this paper hopes to close this gap.
 3 Preliminary
@@
 We study the pool-based active learning problem, where an initial unlabeled set of $N$ examples $X=\{x_{1},...,x_{N}\}$ are available for annotation.
-We let $\mathcal{X}$ denote the space of possible examples.
 Their corresponding labels $Y=\{y_{1},...,y_{N}\}$ are initially unknown.
-Furthermore, we study the multi-class classification problem, so the space of labels $\mathcal{Y}:=[K]$, where each element denotes the class index and $K$ is the total number of classes.
-Moreover, let $N_{1},...,N_{K}$ denote the number of examples in $X$ in each class.
-We define the imbalance ratio as $\gamma=\frac{\min_{k\in[K]}N_{k}}{\max_{k\in[K]}N_{k}}$.
+Furthermore, we study the multi-class classification problem, where the space of labels $\mathcal{Y}:=[K]$ is consisted $K$ classes.
+Moreover, let $N_{1},...,N_{K}$ denote the number of examples in $X$ of each class.
+We define the imbalance ratio as $\gamma=\frac{\min_{k\in[K]}N_{k}}{\max_{k^{\prime}\in[K]}N_{k^{\prime}}}$.
 A deep active learning algorithm iteratively chooses batches of examples for annotation.
 During the $t$-th iteration, the algorithm is given labeled and unlabeled sets of examples, $L_{t}$ and $U_{t}$ respectively, where $L_{t}\cup U_{t}=X$ and $L_{t}\cap U_{t}=\emptyset$.
-The algorithm then chooses $B$ examples from the unlabeled set $X^{(t)}\subseteq U_{t}$ for their corresponding labels $Y^{(t)}$ to be annotated.
+The algorithm then chooses $B$ examples from the unlabeled set $X^{(t)}\subseteq U_{t}$ and then obtains their corresponding labels $Y^{(t)}$.
 The labeled and unlabeled sets are then updated, i.e., $L_{t+1}\leftarrow L_{t}\cup X^{(t)}$ and $U_{t+1}\leftarrow U_{t}\backslash X^{(t)}$.
-Based on new labeled set $L_{t+1}$ and its corresponding labels, a neural network $f_{t}:\mathcal{X}\rightarrow\mathcal{Y}$ is trained to inform the choice for the next iteration.
+Based on new labeled set $L_{t+1}$ and its corresponding labels, a neural network $f_{t}:X\rightarrow[K]$ is trained to inform the choice for the next iteration.
 The ultimate goal of deep active learning is to obtain high predictive accuracy for the trained neural network while annotating as few examples as possible.
@@
 (2022), uncertainty based methods such as confidence (Settles, 2009), margin (Tong & Koller, 2001; Balcan et al., 2006) and entropy (Kremer et al., 2014) sampling could annotate a set of examples with imbalance ratio arbitrarily low.
-The reasoning is simply demonstrated in Figure 0(a), where choosing examples with predictive sigmoid score $\widehat{p}$ closest to .5.5.5.5 may end up annotating examples (all) from the majority class.
+The reasoning is simply demonstrated in Figure 1(a), where choosing examples with predictive sigmoid score $\widehat{p}$ closest to .5.5.5.5 may end up annotating examples almost all from the majority class.
 It is also empirically shown that annotating examples most certain to be in the minority class is not the most effective in improving model performance.
@@
 Annotated examples are shaded.
-Therefore, the key objective in imbalanced active learning is labeling examples that are both uncertain and class balanced.
-The proposed algorithm GALAXY (Zhang et al., 2022) follows the following two-phased procedure: 1.
-Identify the optimal separation threshold $\widehat{p}^{\star}$, as shown in Figure 0(a).
+Therefore, the key objective in imbalanced active learning is to label examples that are both uncertain and class balanced.
+In both GALAXY (Zhang et al., 2022) and this work, we decompose the objective into the following two-phased procedure: 1.
+Identify the optimal separation threshold $j^{\star}$ that best separates the minority class from the majority class, as shown in Figure 1(a).
 2.
-Annotate examples in the minority class closest to $\widehat{p}^{\star}$.
-GALAXY draws inspiration from graph-based active learning in identifying $\widehat{p}^{\star}$.
-In particular, it follows a modified bisection procedure that sequentially annotate examples to search for all cuts, namely thresholds with a class 1 example to the left and a class 2 example to the right (see Figure 0(b)).
-GALAXY then annotate examples close to all such cuts.
-However, the algorithm suffers from two weaknesses.
-Firstly, a bisection procedure only allows sequential labeling, which prevents multiple annotators labeling in parallel.
-Secondly, the problem of identifying the optimal separation threshold is a noisy procedure, since the predictive sigmoid scores may not perfectly separate the minority class examples from the majority.
-GALAXY mitigates the problem by wastefully identifying and annotating around all cuts, which could lead to annotating large number of majority class examples (shown in Figure 0(b)).
+Annotate equal number of examples next to $j^{\star}$ from both sides.
+GALAXY draws inspiration from graph-based active learning in identifying $p^{\star}$.
+In particular, it follows a modified bisection procedure that sequentially annotates examples to find all cuts, namely thresholds with a class 1 example to the left and a class 2 example to the right (see Figure 1(b)).
+As the optimal separation threshold must be located at one of these cuts, GALAXY annotates examples close to all the found cuts.
+However, the algorithm suffers from three weaknesses.
+First, as shown in Figure 1(b), by annotating around all cuts, the algorithm may waste a significant portion of the annotation budget around outliers, which could lead to annotating large number of majority class examples.
+Second, the above problem can be magnified when annotators provide noisy labels.
+Third, a bisection procedure only allows for sequential labeling, which prevents multiple annotators labeling in parallel.
+In this paper, we take a DIRECT approach by identifying only the optimal separation threshold and address all of the shortcomings above.
 4 A Robust Algorithm with Batch Labeling
-In this section, we formally define optimal separation threshold and pose the problem of identifying it as an 1-dimensional reduction to the agnostic active learning problem.
-We then propose a novel algorithm inspired by the agnostic active learning literature (Balcan et al., 2006; Dasgupta et al., 2007; Hanneke et al., 2014; Katz-Samuels et al., 2021).
+In this section, we formally define the optimal separation threshold and pose the problem of identifying it as an 1-dimensional reduction to the agnostic active learning problem.
+We then propose an algorithm inspired by the agnostic active learning literature (Balcan et al., 2006; Dasgupta et al., 2007; Hanneke et al., 2014; Katz-Samuels et al., 2021).
 4.1 An 1-D Reduction to Agnostic Active Learning
 We start by considering the imbalanced binary classification setting mentioned in Section 3.2.
-When given a neural network model, we let $\widehat{p}:\mathcal{X}\rightarrow[0,1]$ be the predictive function mapping examples to sigmoid scores.
+When given a neural network model, we let $\widehat{p}:X\rightarrow[0,1]$ be the predictive function mapping examples to sigmoid scores.
 Here, a higher sigmoid score represents a higher confidence of the example being in class $2$.
-Let $q_{1},...,q_{N}$ denote the sigmoid scores of every example in the pool $X$, where $q_{i}=\widehat{p}(x_{i})$ for each $i\in[N]$.
-We also sort examples by the sigmoid predictive score similar to Section 3.2.
-Formally, we let $q_{(1)}\leq q_{(2)}\leq\cdots q_{(N)}$ be a sorted permutation of $q_{1},...,q_{N}$, and let $y_{(1)},y_{(2)},...,y_{(N)}$ denote the sorted list’s corresponding labels.
-We are now ready to define the optimal separation threshold as described in Section 3.2 below.
+We sort examples by their sigmoid predictive score similar to Section 3.2.
+Formally, we let $q_{(1)}\leq q_{(2)}\leq\cdots\leq q_{(N)}$ be a sorted permutation of $\widehat{p}(x_{1}),...,\widehat{p}(x_{N})$, and let $y_{(1)},y_{(2)},...,y_{(N)}$ denote the sorted list’s corresponding labels.
+We are now ready to define the optimal separation threshold as described in Section 3.2.
 Definition 4.1.
-Given the sorted predictive sigmoid scores $0\leq q_{(1)}\leq\cdots\leq q_{(N)}\leq 1$, we define the optimal separation threshold as $j^{\star}\in[N]$ such that
-$j^{\star}=\operatorname*{arg\,max}_{j}$ (|$\left(\lvert\{y_{(i)}=1:1\leq i\leq j\}\rvert\right.$
-|$\left.-\lvert\{y_{(i)}=2:1\leq i\leq j\}\rvert\right).$ (1)
-In other words, $j^{\star}$ has the largest difference in the number of class $1$ examples from class $2$ examples to its left.
-Furthermore, ties are broken by choosing the largest $j^{\star}$ that attains the argmax if class $1$ is the minority class and the lowest $j^{\star}$ otherwise.
+Given a sorting of $X$ the points according the sigmoid scores $0\leq q_{(1)}\leq\cdots\leq q_{(N)}\leq 1$, we define the optimal separation threshold as $j^{\star}\in[N]$ such that
+$j^{\star}=\operatorname*{arg\,max}_{j}$ (|$\left(\lvert\{y_{(i)}=1:i\leq j\}\rvert-\lvert\{y_{(i)}=2:i\leq j\}\rvert\right)$
+$=\operatorname*{arg\,max}_{j}$ (|$\left(\lvert\{y_{(i)}=2:i>j\}\rvert-\lvert\{y_{(i)}=2:i>j\}\rvert\right).$ (1)
+In other words, on either side of $j^{\ast}$, it has the largest discrepancy in the number of examples between the two classes.
+This captures the intuition of Figure 1(b) — our goal is to find a threshold that best separates one class from the other.
+We quickly remark that ties are broken by choosing the largest $j^{\star}$ that attains the argmax if class $1$ is the minority class and the lowest $j^{\star}$ otherwise.
 1D Reduction.
 We now provide a reduction of finding $j^{\star}$ to an 1-dimensional agnostic active learning problem.
-In particular, let =$\mathcal{H}=\{h_{0},h_{1},...,h_{N}\}$ be the hypothesis class.
-Here each hypothesis $h_{j}:[0,1]\rightarrow\{1,2\}$ maps sigmoid scores to binary target classes and
+We define the hypothesis class =$\mathcal{H}=\{h_{0},h_{1},...,h_{N}\}$ where each hypothesis $h_{j}:[0,1]\rightarrow\{1,2\}$ is defined as
 $h_{j}(q)=\begin{cases}1&\text{if}\,q\leq q_{(j)}\\ 2&\text{if}\,q>q_{(j)}\end{cases}.$
+and $q_{(0)}=0$ defines the hypothesis $h_{0}$ that predicts class $2$ at all times.
 The empirical zero-one loss for each hypothesis is then defined as $\mathcal{L}(h_{j})=\sum_{i=1}^{N}\mathbf{1}\{h_{j}(q_{(i)})\neq y_{(i)}\}$.
-In Appendix Appendix A., we show that optimizing for the zero-one loss $\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$ is equivalent to (1).
+In Appendix A, we show that optimizing for the zero-one loss $\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$ is equivalent to (4.1).
 Namely, $j^{\star}\in\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$, transforming the problem of finding the optimal separation threshold to an 1-D agnostic active learning problem.
@@
 The goal therefore becomes finding all $K$ optimal separation thresholds, which is equivalent with solving $K$ 1-D agnostic active learning problems.
-Moreover, let $\hat{p}:\mathcal{X}\rightarrow\Delta^{(K-1)}$ denote the neural network prediction function mapping examples to sigmoid scores.
-Instead of using sigmoid scores, for class $k$, we use the margin scores $q_{i}^{k}:=\max_{k^{\prime}}[\hat{p}(x_{i})]_{k^{\prime}}-[\hat{p}(x_{i})]_{k}$ to sort the examples.
+Moreover, let $\widehat{p}:X\rightarrow\Delta^{(K-1)}$ denote the neural network prediction function mapping examples to sigmoid scores.
+Instead of using sigmoid scores, for class $k$, we use the margin scores $q_{i}^{k}:=\max_{k^{\prime}}[\widehat{p}(x_{i})]_{k^{\prime}}-[\widehat{p}(x_{i})]_{k}$ to sort the examples.
+When multiple examples have the same margin score, we further break ties by their corresponding confidence scores $[\widehat{p}(x_{i})]_{k}$.
 We also note that sorting with margin scores is equivalent to sorting with sigmoid scores in the binary classification setting, so this is a strictly more general formulation to multi-class classification.
-4.2 Algorithm
-We are now ready to state our algorithm as shown in Algorithm 1.
-The algorithm follows a two-phased, where the first phase aims to identify the optimal separation threshold for each class.
-The second phase then annotates examples closest to the estimated optimal separation thresholds for each class.
   Input: Pool $X$, number of rounds $T$, retraining batch size $B_{\text{train}}$, number of parallel annotations $B_{\text{parallel}}$.
@@
      Initialize labeled set $L_{t}\leftarrow L_{t-1}$.
-     Initialize budget per class $b\leftarrow\frac{B_{\text{train}}}{2K}$.
+     Initialize budget per class $b\leftarrow B_{\text{train}}/2K$.
      for $k$ in $\text{RandPerm}(\{1,...,K\})$ do
-        Compute margin scores $q_{1}^{k},...,q_{N}^{k}$ based on $f_{t-1}$ and sort them by $q_{(1)}^{k}\leq\cdots\leq q_{(N)}^{k}$.
-        Find optimal separation threshold: $L_{t}\leftarrow\text{VReduce}(L_{t},b,k,B_{\text{parallel}},\{(x_{(i)},y_{(i)})\}_{i=1}^{N})$.
+        Compute margin scores $q_{1}^{k},...,q_{N}^{k}$ based on $f_{t-1}$ and sort them by $q_{(1)}^{k}\leq\cdots\leq q_{(N)}^{k}$ and break ties so that $q_{(i)}^{k}=q_{(i+1)}^{k}\implies[\widehat{p}(x_{(i)})]_{k}\geq[\widehat{p}(x_{(i+1)})]_{k}$.
+        Find optimal separation threshold: $L_{t}\leftarrow\text{VReduce}(L_{t},b,k,B_{\text{parallel}},\{q^{k}_{(i)}\}_{i=1}^{N})$.
      end for
      Annotate rare and uncertain examples
-     Compute budget per class $b\leftarrow\frac{B_{\text{train}}-|L_{t}|}{K}$.
+     Compute budget per class $b\leftarrow(B_{\text{train}}-|L_{t}|)/K$.
      for $k$ in $\text{RandPerm}(\{1,...,K\})$ do
@@
 Algorithm 1 DIRECT: DImension REduction for aCTive Learning under Imbalance and Label Noise
-To identify the optimal separation threshold, we loop over each class $k$ and run a fix-budget batch active learning procedure for the one-dimensional reduction (as shown in Algorithm 2).
-Specifically, we keep a version space of the hypotheses with higher likelihood of being the optimal separation threshold.
-We annotate examples sequentially in batches and shrink the version space by a fixed rate each iteration.
-The shrinkage rate $c$ is determined by the budget and batch size, so that after the final iteration, the version space has exactly one hypothesis left.
-To address batch labeling, we note the number of examples one collect before retraining is usually far greater than the number of annotators annotating in parallel.
-In particular, we let $B_{\text{train}}$ denote the number of examples the algorithm collects before the neural network is retrained.
-In practice, this number is usually determined by the practical constraints of computational training cost.
-On the other hand, we let $B_{\text{parallel}}$ denote the number of examples annotated in parallel.
-As discussed above, $B_{\text{parallel}}<<B_{\text{train}}$ in practice.
-Lastly, as will be discussed in Section 6, a variant of our algorithm can also be proposed for asynchronous labeling.
-  Input: Labeled set $L$, budget $b$, class of interest $k$, parallel batch size $B_{\text{parallel}}$, examples and ground truth labels sorted by uncertainty $\{(x_{(i)},y_{(i)})\}_{i=1}^{N}$ (ground truth labels are hidden to the learner).
-  Initialize: Version space: shortest segment of indices $[I,J]$, where $X_{I},X_{J}\in L$, $\forall i\leq I,x_{(i)}\in L\implies y_{(i)}=k$ and $\forall j\geq J,x_{(j)}\in L\implies y_{(j)}\neq k$.
+  Input: Labeled set $L$, budget $b$, class of interest $k$, parallel batch size $B_{\text{parallel}}$, examples and ground truth labels sorted by uncertainty $\{q_{(i)}\}_{i=1}^{N}$.
+  Define: For examples in the labeled set, we let $y_{(i)}$ denote the label of the example corresponding to $q_{(i)}$.
+  Initialize: Version space: shortest segment of indices $[I,J]$, with $q_{(I)},q_{(J)}$ $\forall i\leq I,q_{(i)}\in L\implies y_{(i)}=k$ and $\forall j\geq J,q_{(j)}\in L\implies y_{(j)}\neq k$.
   Initialize: Number of iterations $m\leftarrow\frac{b}{B_{\text{parallel}}}$. Shrinking factor $c\leftarrow\sqrt[m]{J-I}$
   for $t=1,...,m$ do
-     Sample uniformly at random $B_{\text{parallel}}$ unlabeled examples in $x_{(I)}$, …, $x_{(J)}$ for annotation and insert to $L$.
+     Sample uniformly at random $B_{\text{parallel}}$ unlabeled examples in $q_{(I)}$, …, $q_{(J)}$ for annotation and insert to $L$.
      Find $[I^{\prime},J^{\prime}]\subset[I,J]$ such that
 $I^{\prime},J^{\prime}=\operatorname*{arg\,min}_{i,j:j-i=\frac{1}{c}(J-I)}\max\{\widehat{\mathcal{L}}^{k}(i),\widehat{\mathcal{L}}^{k}(j)\}\quad\text{where }$
-^$\widehat{\mathcal{L}}^{k}(s)=\sum_{\begin{subarray}{c}r\leq s:\\ x_{(r)}\in L\end{subarray}}\mathbf{1}\{y_{(r)}\neq k\}+\sum_{\begin{subarray}{c}r>s:\\ x_{(r)}\in L\end{subarray}}\mathbf{1}\{y_{(r)}=k\}.$
+^$\widehat{\mathcal{L}}^{k}(s)=\sum_{\begin{subarray}{c}r\leq s:\\ q_{(r)}\in L\end{subarray}}\mathbf{1}\{y_{(r)}\neq k\}+\sum_{\begin{subarray}{c}r>s:\\ q_{(r)}\in L\end{subarray}}\mathbf{1}\{y_{(r)}=k\}.$
      Update version space $I,J\leftarrow I^{\prime},J^{\prime}$.
@@
 Algorithm 2 VReduce: Version Space Reduction
+4.2 Algorithm
+We are now ready to state our algorithm as shown in Algorithm 1.
+The algorithm follows a two-phased procedure, where the first phase aims to identify the optimal separation threshold for each class.
+The second phase then annotates examples closest to the estimated optimal separation thresholds for each class.
+Our set of labeled examples at anytime are maintained in $L_{t}$.
+To identify the optimal separation threshold, we loop over each class $k$ and run a fix-budget batch active learning procedure for the one-dimensional reduction (as shown in Algorithm 2) on the set of labeled examples.
+Our method is motivated by the fixed-budget algorithm for active classification, ACED (Katz-Samuels et al., 2021).
+Intuitively, in the non-noisy or separable setting, where the loss of $\mathcal{L}(h_{j^{\star}})=0$, we could run a binary search procedure to learn $j^{\star}$.
+However, we can’t assume that $j^{\star}$ perfectly separates the classes.
+ACED circumvents this by maintaining a version space of threshold classifiers, namely an interval $[I,J]$.
+In each of the $b/B_{\text{parallel}}$ rounds, $B_{\text{parallel}}$ samples are taken and appended to the set of labeled examples $L$.
+Then, an interval $I^{\prime},J^{\prime}$ is computed, whose length is a factor of $1/c$ smaller than the current $I,J$ of low-loss classifiers.
+The shrinkage rate $c$ is determined by the budget and batch size, so that after the final iteration, the version space has exactly one hypothesis left.
+The fixed confidence (or alternatively $\delta$-PAC) analogue of ACED is known to match instance optimal lower bounds for active classification up to logarithmic factors.
+Under benign noise conditions, the probability of misidentifying the optimal separation threshold decays exponentially w.r.t.
+budget $b$.
+To address batch labeling, we note the number of examples collected before retraining is usually far greater than the number of annotators annotating in parallel.
+In particular, we let $B_{\text{train}}$ denote the number of examples the algorithm collects before the neural network is retrained.
+In practice, this number is usually determined by the constraints of computational training cost.
+On the other hand, we let $B_{\text{parallel}}$ denote the number of examples annotated in parallel.
+As discussed above, $B_{\text{parallel}}\ll B_{\text{train}}$ in practice.
+Lastly, as will be discussed in Section 6, a variant of our algorithm can also be proposed for asynchronous labeling.
 5 Experiments
-(a) Imbalanced CIFAR-10, two classes
-(b) Imbalanced CIFAR-10, three classes
-(c) Imbalanced CIFAR-100, two classes
-(d) Imbalanced CIFAR-100, three classes
-(e) Imbalanced SVHN, two classes
-(f) Imbalanced SVHN, three classes
-Figure 2: Performance of DIRECT vs different baselines across six different dataset settings.
+We conduct experiments under two primary setups:
+  1. 1.
+    Fine-tuning ResNet-18 with vanilla supervised training on imbalanced datasets similar to Zhang et al. (2022).
+  2. 2.
+    Fine-tuning large pretrained model (CLIP ViT-B32) with semi-supervised training strategies under the LabelBench framework (Zhang et al., 2023a).
+In Section 5.1, we document these setups in detail.
+For both evaluation setups, we first evaluate the performance of DIRECT under the noiseless setting in Section 5.2, showing its superior label-efficiency and ability to accommodate batch labeling.
+In Section 5.3, we evaluate deep active learning algorithms under a novel setting with both class imbalance and noisy labels.
+Under this setting, we also include an ablation study of the performance of DIRECT on various levels of label noises.
+While we highlight many results in this section, see Appendix B for all results.
 5.1 Experiment Setup
-Our experiment setup follows closely from Zhang et al.
-(2022).
-We use the CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009) and SVHN (Netzer et al., 2011) image classification datasets.
-The original CIFAR-10 and SVHN datasets have 10 classes while the CIFAR-100 dataset has 100 classes.
+ResNet-18 Experiments.
+ResNet-18 with passive training has been the standard evaluation in existing deep active literature (Ash et al., 2019; Zhang et al., 2022).
+Our experiment setup utilizes the CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), SVHN (Netzer et al., 2011) and PathMNIST (Yang et al., 2021) image classification datasets.
+The original forms of these datasets are roughly balanced across 9, 10 or 100 classes.
 We construct an extremely imbalanced dataset by grouping a large number of classes into one majority class.
-For example, suppose we have a dataset with $M$ classes.
-We generate an imbalanced dataset with $K$ classes ($K<M$) by reusing classes $11,...K-$ 1 from the original dataset and combining the rest of the classes $K,...,M$ into a single majority class $K$.
-Detailed imbalance ratios are shown in Table 1.
-Name $K$ $N_{K}$ $\sum_{k=1}^{K-1}N_{k}$ Imb Ratio $\gamma$
-CIFAR-10 $2$ $45000$ $5000$$1$.1111.1111.1111
-CIFAR-10 $3$ $40000$ $10000$1250.1250.1250.1250
-CIFAR-100 $2$ $49500$ $500$0101.0101.0101.0101
-CIFAR-100 $3$ $49000$ $1000$0102.0102.0102.0102
-SVHN $2$ $68309$ $4948$0724.0724.0724.0724
-SVHN $3$ $54448$ $18809$2546.2546.2546.2546
-Table 1: Dataset setting for extremely unbalanced scenario.
-$N_{K}$ denotes the number of images in the majority class while $\sum_{k=1}^{K-1}N_{k}$ is the total number of images in all minority classes.
-$\epsilon$ is the class imbalance factor defined in Section 3.1.
-We compare our algorithm DIRECT against nine baselines: GALAXY (Zhang et al., 2022), SIMILAR (Kothawade et al., 2021), BADGE (Ash et al., 2019), BASE (Emam et al., 2021), BAIT (Ash et al., 2021), Cluster Margin (Citovsky et al., 2021), Confidence Sampling (Settles, 2009), Most Likely Positive (Jiang et al., 2018; Warmuth et al., 2001, 2003) and Random Sampling.
-The highlight surrounding the baselines represents the standard error plotted as the confidence intervals.
-Due to computational constraints, we only have single runs for SIMILAR, BADGE, and BAIT.
-For all other baselines, we average over four runs.
-For all experiments, we use PyTorch’s ResNet-18 model pretrained on ImageNet for initialization. We use the Adam optimization algorithm with 500 epochs for each L and the learning rate set to 0.01.
-5.2 Results
-Synchronous Labeling ($B_{\text{parallel}}=1$).
-In our first set of experiments, we compare DIRECT against GALAXY in the synchronous labeling setting by setting $B_{\text{parallel}}=1$ in our algorithm.
-As shown in Figure 2(a), DIRECT outperforms GALAXY by saving an additional $15\%$ of the labeling budget.
-Compared to random sampling, DIRECT is able to around $90\%$ of annotation cost.
-Batch Labeling.
-We observe that on all six datasets shown in Figure 2, DIRECT performs comparably to GALAXY even when under parallel annotation ($B_{\text{parallel}}=5$).
-However, we do observe a slight drop in label-efficiency when comparing to synchronous labeling.
-6 Future Work
-In this paper, we addressed the batch sampling problem of GALAXY by annotating multiple examples in parallel.
-However, in practice, examples are often annotated asynchronously.
-We believe this is a valuable future direction and a potential solution is to utilize an asynchronous variant of one-dimensional active learning algorithm.
+For example, given a balanced dataset above with $M$ classes.
+We generate an imbalanced dataset with $K$ classes ($K<M$) by the first $K-1$ classes from the original dataset and combining the rest of the classes $K,...,M$ into a single majority class $K$.
+Imbalance ratios are shown in Table 1.
+For neural network training, we utilize the standard passive training on labeled examples with cross entropy loss and Adam optimizer (Kingma & Ba, 2014).
+The ResNet-18 model (He et al., 2016) is pretrained on ImageNet (Deng et al., 2009) from the PyTorch library.
+To address data imbalance, we utilize a reweighted cross entropy loss by the inverse frequency of the number of labeled examples in each class.
+For experiments with label noise, we further add a $10\%$ label smoothing during training (Müller et al., 2019).
+LabelBench Experiments.
+Proposed by Zhang et al.
+(2023a), LabelBench evaluates active learning performance in a more comprehensive framework.
+Here, we fine-tune the large pretrained model from CLIP’s ViT-B32 model (Radford et al., 2021).
+The framework also utilizes semi-supervised learning method FlexMatch (Zhang et al., 2021) to further leverage the unlabeled examples in the pool for training.
+We conduct experiments on the two imbalanced datasets in LabelBench, with FMoW (Christie et al., 2018) and iWildcam (Beery et al., 2021).
+Similar to the ResNet-18 experiments, we use a $10\%$ label smoothing in the loss function to improve training under label noise.
+We did find FlexMatch to perform poorly under the combination of imbalance and label noise, so we used the passive training method for label noise experiments.
+Name # Classes $K$ $N$ Imb Ratio $\gamma=\frac{\min_{k}N_{k}}{\max_{k^{\prime}}N_{k^{\prime}}}$
+Imb CIFAR-10 $2$ $50000$$1$.1111.1111.1111
+Imb CIFAR-10 $3$ $50000$1250.1250.1250.1250
+Imb CIFAR-100 $2$ $50000$0101.0101.0101.0101
+Imb CIFAR-100 $3$ $50000$0102.0102.0102.0102
+Imb CIFAR-100 $10$ $50000$0110.0110.0110.0110
+Imb SVHN $2$ $73257$0724.0724.0724.0724
+Imb SVHN $3$ $54448$2546.2546.2546.2546
+PathMNIST $2$ $89996$1166.1166.1166.1166
+FMoW $62$ $76863$0049.0049.0049.0049
+iWildCam $14$ $129809$ $4.57\cdot 10^{-5}$
+Table 1: Dataset settings for our experiments. $N$ denotes the total number of examples in our dataset. $\gamma$ is the class imbalance ratio defined in Section 3.1.
+(a) Imbalanced CIFAR-10, three classes
+(b) Imbalanced CIFAR-100, ten classes
+(c) Imbalanced SVHN, two classes
+(d) PathMNIST, two classes
+(e) FMoW, Balanced Pool Accuracy
+(f) Imbalanced SVHN, two classes, number of labeled examples in minority class
+Figure 3: Performance comparison of DIRECT against other baselines algorithms in the noiseless but imbalanced setting.
+(a)-(d) are balanced accuracy on ResNet-18 experiments while (e) shows experiment under the LabelBench framework.
+$B_{\text{parallel}}$ indicates the number of parallel annotations as mentioned in Section 4.2.
+$B_{\text{parallel}}=1$ is equivalent with sychornous labeling.
+Results are averaged over four trials and the shaded areas represent standard errors around the average performances.
+5.2 Experiments under Imbalance, without Label Noise
+For the noiseless experiment on ResNet-18, we compare against nine baselines: GALAXY (Zhang et al., 2022), SIMILAR (Kothawade et al., 2021), BADGE (Ash et al., 2019), BASE (Emam et al., 2021), BAIT (Ash et al., 2021), Cluster Margin (Citovsky et al., 2021), Confidence Sampling (Settles, 2009), Most Likely Positive (Jiang et al., 2018; Warmuth et al., 2001, 2003) and Random Sampling.
+We briefly distinguish the algorithms into two categories.
+In particular, both SIMILAR and Most Likely Positive annotate examples that are similar to existing labeled minority examples, thus can significantly annotate a large quantity of minority examples.
+The rest of the algorithms primarily optimizes for different notions of informativeness such as diversity and uncertainty.
+For the LabelBench experiments, due to the large dataset and model embedding sizes, we choose a subset of the algorithms that are computationally efficient and among top performers in the ResNet-18 results.
+In particular, we compared against BADGE, Margin Sampling (Balcan et al., 2006; Kremer et al., 2014), CORESET and GALAXY.
+As highlighted in Figures 1(a) and 3(a)-(d), DIRECT consistently and significantly outperforms existing algorithms on the ResNet-18 experiments.
+In Figures 1(c) and 3(e), we demonstrate the increased label-efficiency is also consistently shown in the LabelBench experiments.
+Compared to random sampling, DIRECT can save more than 80% of the annotation cost on imbalanced SVHN experiment of Figure 3(c).
+In terms of class-balancedness, we consistently observe that both Most Likely Positive and SIMILAR annotating greater number of minority class examples, but significantly underperforms in terms of balanced accuracy (an example showin in Figure 3(f)).
+While Zhang et al.
+(2022) has already observed this phenomenon, we can further see that DIRECT collects slightly less minority class examples than GALAXY, but outperforms in terms of balanced accuracy.
+While it is crucial to optimize class-balancedness for better model performance, we see that both extremes of annotating too few and too many minority examples could lead to worse generalization performances.
+When too few examples are from minority class, the performance of the minority classes could be significantly hindered.
+When optimized to annotate as many examples from minority class as possible, the algorithm has to tradeoff annotating informative examples to examples it is more certain to be in the minority class.
+Together, this suggests an intricate balance between the two objectives, generalization performance and class-balancedness.
+We would also like to highlight the ability to handle batch labeling.
+Across our experiments, we see DIRECT outperforms with different amounts of parallel annotation ($B_{\text{parallel}}=1,5$ and $20$), indicating its general effectiveness.
+This is in comparison to the synchoronous nature of GALAXY, where it is always using $B_{\text{parallel}}=1$.
+On Figures 1(a) and 4(a), we see that DIRECT outperforms GALAXY with synchronous labeling.
+Furthermore, in these experiments we also see using $B_{\text{parallel}}=5$ only affects algorithm performances minimally for DIRECT.
+(a) Imbalanced CIFAR-10, two classes, 10% label noise
+(b) Imbalanced SVHN, two classes, 10% label noise
+(c) iWildcam Balanced Pool Accuracy, $10\%$ label noise
+Figure 4: Performance of DIRECT against baseline algorithms under $10\%$ label noise.
+(a)-(b) are balanced accuracy on ResNet-18 experiments while (c) shows results under the LabelBench framework.
+Results are averaged over four trials and the shaded areas represent standard errors around the average performances.
+(a) Imbalanced CIFAR-100, two classes, no label noise
+(b) Imbalanced CIFAR-100, two classes, 10% label noise
+(c) Imbalanced CIFAR-100, two classes, 15% label noise
+Figure 5: Performance of DIRECT against baseline algorithms under different levels of label noise.
+Results are averaged over four trials and the shaded areas represent standard errors around the average performances.
+5.3 Experiments under Imbalance and Label Noise
+We conduct novel sets of experiments under both class imbalance and label noise.
+Here, for both ResNet-18 and LabelBench experiments, we evaluate against all of the algorithms that performed well under the imbalance but noiseless setting above.
+For all of our experiments, we introduce a fixed percentage of label noise, where the given fraction of the examples’ labels are corrupted to a different class uniformly at random.
+For most of our experiements with $10\%$ label noise shown in Figure 4, we observe again that DIRECT consistently improves over all baselines including GALAXY.
+The results are consistent on ResNet-18 and LabelBench setups, and with different $B_{\text{parallel}}$ values, showing DIRECT’s robustness under label noise.
+Different Levels of Label Noise As shown in Figures 5(a)-(c) and 1(b), we observe the results on imbalanced CIFAR-100 with two classes across numerous levels of label noise, with $0\%$, $10\%$, $15\%$ and $20\%$ respectively.
+In fact, the noiseless experiment in Figure 5(a) is the only setting DIRECT slightly underperforms GALAXY in terms of generalization accuracy.
+However, we see DIRECT becomes more label-efficient under the more practical label noise settings.
+It is also worth noting that with high label noise of $20\%$, we observe in Figure 1(b) that existing algorithms underperform random sampling.
+In contrast, DIRECT significantly outperforms random sampling, saving more than $60\%$ of the annotation cost.
+6 Conclusion and Future Work
+In this paper, we conducted the first study of deep active learning under both class imbalance and label noise.
+We proposed an algorithm DIRECT that significantly and consistently outperforms existing literature.
+In this work, we also addressed the batch sampling problem of current SOTA algorithm, GALAXY (Zhang et al., 2022), by annotating multiple examples in parallel.
+Studying asynchronous labeling could a natural extension of our work.
+We think a potential solution is to utilize an asynchronous variant of one-dimensional active learning algorithm.
+In addition, one can further batch the labeling process across different classes to further accommodate an even larger number of parallel annotators.
 Acknowledgements
 This work has been supported in part by NSF Award 2112471.
-Appendix A.
-Lemma .1.
+Impact Statement
+In the rapidly evolving landscape of machine learning, the efficacy of active learning in addressing data imbalance and label noise is a significant stride towards more robust and equitable AI systems.
+This research explores how active learning can effectively mitigate the challenges posed by imbalanced datasets and erroneous labels, prevalent in real-world scenarios.
+The positive impacts of this research are multifaceted.
+It enhances the accessibility and utility of machine learning in domains where data imbalance is a common challenge, such as healthcare, finance, and social media analytics.
+By improving class-balancedness in annotated sets, models trained on these datasets are less biased and more representative of real-world distributions, leading to fairer and more accurate outcomes.
+Additionally, this research contributes to reducing the time and cost associated with data annotation, which is particularly beneficial in fields where expert annotation is expensive or scarce.
+However, if not carefully implemented, active learning strategies could inadvertently introduce new biases or amplify existing ones, particularly in scenarios where the initial data is severely imbalanced or contains deeply ingrained biases.
+Furthermore, the advanced nature of these techniques may widen the gap between organizations with access to state-of-the-art technology and those without, potentially exacerbating existing inequalities in technology deployment.
+Appendix A Equivalent Objective
+Lemma A.1.
 The agnostic active learning reduction is equivalently finding the optimal separation threshold. Namely,
@@
 $=\operatorname*{arg\,max}_{j}\sum_{i=1}^{j}\left(\mathbf{1}\{y_{(i)}=1\}-\mathbf{1}\{y_{(i)}=2\}\right)$
-lt="[LOGO]">
+Appendix B All Results
+B.1 Noiseless Results under Imbalance
+Figure 6: Imbalanced CIFAR-10, two classes.
+Figure 7: Imbalanced CIFAR-10, three classes.
+Figure 8: Imbalanced CIFAR-100, two classes.
+Figure 9: Imbalanced CIFAR-100, three classes.
+Figure 10: Imbalanced CIFAR-100, 10 classes.
+Figure 11: Imbalanced SVHN, two classes.
+Figure 12: Imbalanced SVHN, three classes.
+Figure 13: PathMNIST, two classes.
+(a) FMoW Balanced Pool Accuracy
+(b) iWildcam Balanced Pool Accuracy
+Figure 14: LabelBench results in the noiseless setting.
+B.2 Label Noise Results under Imbalance
+Figure 15: Imbalanced CIFAR-10, two classes, 10% label noise.
+Figure 16: Imbalanced CIFAR-10, three classes, 10% label noise.
+Figure 17: Imbalanced CIFAR-100, two classes, 10% label noise.
+Figure 18: Imbalanced CIFAR-100, two classes, 15% label noise.
+Figure 19: Imbalanced CIFAR-100, two classes, 20% label noise.
+Figure 20: Imbalanced CIFAR-100, three classes, 10% label noise.
+Figure 21: Imbalanced SVHN, two classes, 10% label noise.
+Figure 22: Imbalanced SVHN, three classes, 10% label noise.
+(a) FMoW Balanced Pool Accuracy
+(b) iWildcam Balanced Pool Accuracy
+Figure 23: LabelBench results in the 10% label noise setting.

### v3

@@
 We propose a novel algorithm that robustly identifies the class separation threshold and annotates the most uncertain examples that are closest from it.
-Through a novel reduction to one-dimensional active learning, our algorithm DIRECT is able to leverage the classic active learning literature to address issues such as batch labeling and tolerance towards label noise.
+Through a novel reduction to one-dimensional active learning, our algorithm DIRECT is able to leverage classic active learning theory and methods to address issues such as batch labeling and tolerance towards label noise.
 We present extensive experiments on imbalanced datasets with and without label noise.
 Our results demonstrate that DIRECT can save more than 60% of the annotation budget compared to state-of-art active learning algorithms and more than 80% of annotation budget compared to random sampling.
-Machine Learning, ICML
 1 Introduction
 Large-scale deep learning models are playing increasingly important roles across many industries.
-Human feedback and annotations have played a significant role in developing such systems through.
-Progressively over time, we believe the role of humans in a machine learning pipeline would shift to annotating rare yet important cases.
+Human feedback and annotations have played a significant role in developing such systems.
+Progressively over time, we believe the role of humans in a machine learning pipeline will shift to annotating rare yet important cases.
 However, under data imbalance, the typical strategy of randomly choosing examples for annotation becomes especially inefficient.
 This is because the majority of the labeling budget would be spent on common and well-learned classes, resulting in insufficient rare class examples for training an effective model.
-For many annotation jobs, this challenge of data imbalance is further compounded by label noises induced from annotators.
-In this paper, we mitigate these issues by proposing a deep active learning strategy.
+To mitigate this issue, many recent active learning algorithms have focused on labeling more class-balanced and informative examples (Aggarwal et al., 2020; Kothawade et al., 2021; Zhang et al., 2022, 2024b; Soltani et al., 2024).
+For many large-scale annotation jobs, this challenge of data imbalance is further compounded by label noise – a critical and common issue that results from annotator decision fatigue and perception differences.
+A rich body of literature on agnostic active learning (Balcan et al., 2006; Dasgupta et al., 2007; Hanneke et al., 2014; Katz-Samuels et al., 2021) addresses this challenge on low-complexity model classes (e.g.
+linear models).
+However, for deep learning models, these algorithms often becomes ineffective due to the large model class complexity.
+In this paper, we propose a novel active learning strategy for both class imbalance and label noise.
 Our algorithm DIRECT sequentially and adaptively chooses informative and more class-balanced examples for annotation while being robust to noisy annotations.
 To the best of our knowledge, this is the first deep active learning study to address the challenging yet prevalent scenario where both imbalance and label noise coexist.
-Inspired by previous work by Zhang et al.
-(2022), we propose a novel reduction of the imbalanced classification problem into one-dimensional active learning problems.
+To bridge the gap between the imbalanced deep active learning and the agnostic active learning literature, we propose a novel reduction of the imbalanced classification problem into a set of one-dimensional agnostic active learning problems.
 For each class, our reduction sorts unlabeled examples into an list ordered by one-vs-rest margin scores.
-The goal of DIRECT is to find the optimal separation threshold which best separates the examples in any given class from the rests.
-With an equivalence to finding the best threshold classifier, we were able to employ classic active learning literature for obtaining the separation threshold.
-Once the optimal separation threshold is identified, DIRECT annotates a set of more class-balanced examples near this threshold.
-As examples around the optimal separation threshold tend to have high uncertainty, DIRECT annotates examples that are both class-balanced and informative.
-With the reduction to classic one-dimensional active learning literature, DIRECT also inherits the advantages from classic active learning algorithms such as robustness against label noise and batch labeling.
-This is in contrast to the state-of-art imbalanced active learning algorithm GALAXY proposed by Zhang et al.
-(2022), which requires sequential labeling and is much less robust against label noise.
-As demonstrated in Figure 1, under both noisy and noiseless label settings, DIRECT improves significantly over GALAXY’s label efficiency – DIRECT requires less annotations to achieve the same accuracy.
-It achieves state-of-art label efficiency gain on both supervised fine-tuning of ResNet-18 and semi-supervised fine-tuning of large pretrained model under the LabelBench framework (Zhang et al., 2023a).
-Lastly, DIRECT also allows for parallel annotations with various batch sizes.
+The objective of DIRECT is to find the optimal separation threshold which best separates the examples in the given class from the rest.
+By relating our problem to that of finding the best threshold classifier, we are able to employ ideas from the agnostic active learning literature to learn the separation threshold robustly under label noise.
+By annotating around the threshold, the annotated examples are more class-balanced and informative.
+Comparing to existing active learning algorithms such as BADGE (Ash et al., 2019), Cluster-Margin (Citovsky et al., 2021), SIMILAR (Kothawade et al., 2021), GALAXY (Zhang et al., 2022) and many others, DIRECT improves significantly in label efficiencies – less annotations needed to reach the same accuracy.
+Notably, most existing methods mentioned above are proposed to handle batch labeling, while previous work by Zhang et al.
+(2022) proposes a superior performance algorithm at the cost of only allowing one annotation at a time.
+Our algorithm DIRECT is able to obtain the best of both worlds – practical scalability to large annotation jobs by batch labeling while also getting superior performance than all algorithms including GALAXY.
+On imbalanced datasets, DIRECT achieves state-of-art label efficiency on both supervised fine-tuning of ResNet-18 and semi-supervised fine-tuning of large pretrained model under the LabelBench (Zhang et al., 2024a) framework.
 (a) Imbalanced CIFAR-10, two classes, no label noise.
@@
 Both (a) and (b) are using supervised training of ResNet-18.
-In (c), we finetune CLIP ViT-B32 model in combination of semi-supervised training under the LabelBench framework (Zhang et al., 2023a).
-$B_{\text{parallel}}$ is the batch size indicating the number of parallel annotators for the annotation job.
+In (c), we finetune CLIP ViT-B32 model in combination of semi-supervised training under the LabelBench framework (Zhang et al., 2024a).
+$B_{\text{parallel}}$ is the batch size indicating the number of parallel annotators.
 $B_{\text{parallel}}=1$ indicates the synchronous annotation requirement by GALAXY.
-Our algorithm DIRECT takes pre-specified $B_{\text{parallel}}$ as input, which is determined by real world decisions.
+Our algorithm DIRECT takes pre-specified $B_{\text{parallel}}$ as input, which is determined by real world scenarios.
 To summarize our main contributions:
-  • We propose a novel reduction that bridges the advancement in theoretical active learning literature to imbalanced active classification for deep neural networks.
+  • We propose a novel reduction that bridges the advancement in the theoretical agnostic active learning literature to imbalanced active classification for deep neural networks.
   • Our algorithm DIRECT addresses the prevalent imbalance and label noise issues and annotates a more class-balanced and informative set of examples.
@@
   • We conduct experiments across eight dataset settings, four levels of label noise and for both ResNet-18 and large pretrained model (CLIP ViT-B32).
-DIRECT consistently performs the best compared against existing baseline algorithms.
-Quantitatively, DIRECT can save more than $60\%$ annotation cost compared to GALAXY, and more than $80\%$ annotation cost compared to random sampling.
+DIRECT consistently outperforms existing baseline algorithms by saving more than $60\%$ annotation cost compared to the best existing algorithm, and more than $80\%$ annotation cost compared to random sampling.
 2 Related Work
-Class-Balanced Deep Active Learning Active Learning strategies sequentially and adaptively choose examples for annotation (Settles, 2009).
-Many uncertainty-based deep active learning methods extend the traditional active learning literature such as margin, least confidence and entropy sampling (Tong & Koller, 2001; Settles, 2009; Balcan et al., 2006; Kremer et al., 2014).
-These methods have also been shown to perform among top methods when fine-tuning large pretrained models and combined with semi-supervised learning algorithms (Zhang et al., 2023a).
-More sophisticated methods have also been proposed to optimize the chosen unlabeled examples’ uncertainty (Gal et al., 2017; Ducoffe & Precioso, 2018; Beluch et al., 2018), diversity (Sener & Savarese, 2017; Geifman & El-Yaniv, 2017; Citovsky et al., 2021), or a mix of both (Ash et al., 2019, 2021; Wang et al., 2021; Elenter et al., 2022; Mohamadi et al., 2022).
-However, these methods often perform poorly under prevalent and realistic scenarios such as under label noises (Khosla et al., 2022) or under class imbalance (Kothawade et al., 2021; Zhang et al., 2022, 2023a).
+Class-Balanced Deep Active Learning Active earning strategies sequentially and adaptively choose examples for annotation.
+Many uncertainty-based deep active learning methods extend the traditional active learning literature such as margin, least confidence and entropy sampling (Tong and Koller, 2001; Settles, 2009; Balcan et al., 2006; Kremer et al., 2014).
+These methods have been shown to perform among the top when fine-tuning large pretrained models and combined with semi-supervised learning algorithms (Zhang et al., 2024a).
+More sophisticated methods have been proposed to optimize chosen examples’ uncertainty (Gal et al., 2017; Ducoffe and Precioso, 2018; Beluch et al., 2018), diversity (Sener and Savarese, 2017; Geifman and El-Yaniv, 2017; Citovsky et al., 2021), or a mix of both (Ash et al., 2019, 2021; Wang et al., 2021; Elenter et al., 2022; Mohamadi et al., 2022).
+However, these methods often perform poorly under prevalent and realistic scenarios such as label noises (Khosla et al., 2022) or class imbalance Kothawade et al.
+(2021); Zhang et al.
+(2022, 2024a).
 Deep Active Learning under Imbalance Data imbalance and rare instances are prevalent in almost all modern machine learning applications.
-Active learning techniques are effective in addressing the problem in its root by collecting a more class-balanced dataset (Aggarwal et al., 2020; Kothawade et al., 2021; Emam et al., 2021; Zhang et al., 2022; Coleman et al., 2022; Jin et al., 2022; Cai, 2022; Zhang et al., 2023b).
+Active learning techniques are effective in addressing the problem in its root by collecting a more class-balanced dataset (Aggarwal et al., 2020; Kothawade et al., 2021; Emam et al., 2021; Zhang et al., 2022; Coleman et al., 2022; Jin et al., 2022; Cai, 2022; Zhang et al., 2024b).
 To this end, Kothawade et al.
 (2021) propose a submodular-based method that actively annotates examples similar to known examples of rare instances.
-Our work is motivated by Zhang et al.
-(2022), where they propose GALAXY, an algorithm that construct one-dimensional linear graphs and apply graph-based active learning techniques in annotating a set of examples that are both class-balanced and uncertain.
+GALAXY(Zhang et al., 2022) constructs one-dimensional linear graphs and applies graph-based active learning techniques in annotating a set of examples that are both class-balanced and uncertain.
 While GALAXY outperforms existing algorithms, due to a bisection procedure involved, it does not allow parallel annotation.
@@
 Unlike our goal of proposing a new deep active learning algorithm, the paper proposes meta algorithms to choose the right active learning algorithm among a large number of candidate algorithms.
+Agnostic Active Learning for Label Noise Label noise for active learning has been primarily studied under the extensive literature on agnostic learning.
+We refer the interested reader to the survey (Hanneke et al., 2014) for a thorough discussion.
+All of these works, beginning with the seminal works by Balcan et al.
+(2006); Dasgupta et al.
+(2007), follow a familiar paradigm of disagreement based learning.
+This involves maintaining a version space of promising hypotheses at each time and constructing a disagreement region of unlabeled examples.
+For any unlabeled example in the disagreement region, there exists two hypotheses in the version disagreeing on their predictions.
+An example then chosen for annotation by sampling from a informative sampling distribution computed over the disagreement region.
+Several approaches have been proposed for computing such sampling distributions, e.g.
+Jain and Jamieson (2019); Katz-Samuels et al.
+(2020, 2021); Huang et al.
+(2015).
+As described in Section 4.2, our main subroutine VReduce is equivalent to fixed-budget one dimensional threshold disagreement learning based on the ACED algorithm of Katz-Samuels et al.
+(2021).
+We remark that these algorithms tend to be overly pessimistic in training deep neural nets, and this paper hopes to close this gap.
 Deep Active Learning under Label Noise Label noisy settings has rarely been studied in the deep active learning literature.
@@
 This is fundamentally different from our work, where our annotator may provide noisy labels.
-Another line of more theoretical active learning research studies active learning with multiple annotators with different qualities (Zhang & Chaudhuri, 2015; Chen et al., 2022).
-The primary goal in these work is to identify examples a weak annotator and a strong annotator may disagree, in order to only use the strong annotator on those instances.
+Another line of more theoretical active learning research studies active learning with multiple annotators with different qualities (Zhang and Chaudhuri, 2015; Chen et al., 2022).
+The primary goal in these work is to identify examples a weak annotator and a strong annotator may disagree, in order to only use the strong annotator on such instances.
 In our work, we assume access to a single source of annotator that is noisy, which is prevalent in annotation jobs today.
@@
 Unlike their work, our work is agnostic to the noise distributions and conduct experiments on uniformly random corrupted labels.
-To the best of our knowledge, no existing deep active learning work has studied the scenario where both imbalance and label noise present.
+To our knowledge, no deep active learning literature has studied the scenario where both imbalance and label noise present.
 Yet, this setting is the most prevalent in real-world annotation applications.
-Agnostic Active Learning A critical part of our solution is the reduction to a one dimensional agnostic active learning problem, e.g.
-Section 4.1.
-There is extensive work on agnostic learning in general, and we refer the interested reader to the survey (Hanneke et al., 2014) for a thorough discussion.
-All of these works, beginning with the seminal works (Balcan et al., 2006; Dasgupta et al., 2007) follow, a familiar paradigm of disagreement based learning.
-This involves maintainin a version space of promising hypothesis at each time, and constructing the disagreement region as a set of unlabeled examples for which there exists two hypotheses that disagree on their predictions.
-The example to label is chosen by sampling from a informative sampling distribution computed over the disagreement region.
-Several approaches have been proposed for computing the sampling distribution, e.g.
-Jain & Jamieson (2019); Katz-Samuels et al.
-(2020, 2021); Huang et al.
-(2015).
-As described in Section 4.2, our main subroutine VReduce is equivalent to fixed-budget one dimensional threshold disagreement learning based on the ACED algorithm of Katz-Samuels et al.
-(2021).
-We remark that the problem of class imbalance has been largely ignored by the classical work on traditional agnostic learning, and this paper hopes to close this gap.
 3 Preliminary
@@
 Their corresponding labels $Y=\{y_{1},...,y_{N}\}$ are initially unknown.
-Furthermore, we study the multi-class classification problem, where the space of labels $\mathcal{Y}:=[K]$ is consisted $K$ classes.
+Furthermore, we study the multi-class classification problem, where the space of labels $\mathcal{Y}:=[K]$ is consisted of $K$ classes.
 Moreover, let $N_{1},...,N_{K}$ denote the number of examples in $X$ of each class.
@@
 The ultimate goal of deep active learning is to obtain high predictive accuracy for the trained neural network while annotating as few examples as possible.
-3.2 Review of Imbalanced Active Learning: GALAXY
-We first consider a binary case of imbalanced classification, where we assume $N_{1}<N_{2}$ without loss of generality.
-After a total of $T$ rounds and sampling $TB$ examples, random sampling would annotate a subset of $X$ with an imbalance ratio close to $\frac{N_{1}}{N_{2}}$.
-On the other hand, as shown previously in Zhang et al.
-(2022), uncertainty based methods such as confidence (Settles, 2009), margin (Tong & Koller, 2001; Balcan et al., 2006) and entropy (Kremer et al., 2014) sampling could annotate a set of examples with imbalance ratio arbitrarily low.
-The reasoning is simply demonstrated in Figure 1(a), where choosing examples with predictive sigmoid score $\widehat{p}$ closest to .5.5.5.5 may end up annotating examples almost all from the majority class.
-It is also empirically shown that annotating examples most certain to be in the minority class is not the most effective in improving model performance.
+3.2 Limitations of Existing Imbalanced Active Learning Algorithms
 […]
 (b) GALAXY spends approximately equal annotation budget around both cuts, while the cut on the right would yield examples mostly in the majority class.
-Figure 2: Figures demonstrating GALAXY.
+Figure 2: Demonstration of existing imbalance active learning algorithms.
 Ordered lists of examples are ranked by the predictive sigmoid score $\widehat{p}$.
-The ground truth label of each example is represented by its border – solid blue lines for class $1$ and dotted red lines for class $2$.
+The ground truth label of each example is represented by its border – solid blue for class $1$ and dotted red for class $2$.
 Annotated examples are shaded.
-Therefore, the key objective in imbalanced active learning is to label examples that are both uncertain and class balanced.
-In both GALAXY (Zhang et al., 2022) and this work, we decompose the objective into the following two-phased procedure: 1.
-Identify the optimal separation threshold $j^{\star}$ that best separates the minority class from the majority class, as shown in Figure 1(a).
-2.
-Annotate equal number of examples next to $j^{\star}$ from both sides.
-GALAXY draws inspiration from graph-based active learning in identifying $p^{\star}$.
-In particular, it follows a modified bisection procedure that sequentially annotates examples to find all cuts, namely thresholds with a class 1 example to the left and a class 2 example to the right (see Figure 1(b)).
-As the optimal separation threshold must be located at one of these cuts, GALAXY annotates examples close to all the found cuts.
-However, the algorithm suffers from three weaknesses.
-First, as shown in Figure 1(b), by annotating around all cuts, the algorithm may waste a significant portion of the annotation budget around outliers, which could lead to annotating large number of majority class examples.
-Second, the above problem can be magnified when annotators provide noisy labels.
-Third, a bisection procedure only allows for sequential labeling, which prevents multiple annotators labeling in parallel.
+Below we document the several active learning algorithms and how their progressive improvement.
+At the end, we highlight the shortcomings of the state-of-art algorithm GALAXY (Zhang et al., 2022) and motivate DIRECT’s objective of adaptively finding the optimal separation threshold.
+We first consider an imbalanced binary classification case, where $N_{1}<N_{2}$ without loss of generality.
+Random Sampling.
+After annotating a significant number of examples, random sampling would annotate a subset of $X$ with an imbalance ratio close to $\frac{N_{1}}{N_{2}}$.
+This approach suffers from annotating examples that are neither class-balanced nor informative.
+[…]
+[…]
+[…]
+  1. 1.
+    During active learning, the neural network is still under training and cannot perfectly separate the two classes of examples yet.
+Therefore, the sorted graph could have a significant number of cuts.
+As an example in Figure 2(b), when annotating around all of such cuts, the algorithm could waste a significant portion of the annotation budget around misclassified outliers, leading to a large number of majority class annotations.
+  2. 2.
+    Under label noise, the incorrect annotation could lead to more cuts in the sorted graph, further exacerbating the above issue.
+  3. 3.
+    GALAXY finds all cuts through a modified bisection procedure, which only allows for sequential labeling and prevents multiple annotators labeling in parallel.
 In this paper, we take a DIRECT approach by identifying only the optimal separation threshold and address all of the shortcomings above.
-4 A Robust Algorithm with Batch Labeling
+4 A Robust Algorithm for Active Learning under Imbalance and Label Noise
 In this section, we formally define the optimal separation threshold and pose the problem of identifying it as an 1-dimensional reduction to the agnostic active learning problem.
@@
 We sort examples by their sigmoid predictive score similar to Section 3.2.
-Formally, we let $q_{(1)}\leq q_{(2)}\leq\cdots\leq q_{(N)}$ be a sorted permutation of $\widehat{p}(x_{1}),...,\widehat{p}(x_{N})$, and let $y_{(1)},y_{(2)},...,y_{(N)}$ denote the sorted list’s corresponding labels.
-We are now ready to define the optimal separation threshold as described in Section 3.2.
+Formally, we now define the optimal separation threshold as described in Section 3.2.
 Definition 4.1.
-Given a sorting of $X$ the points according the sigmoid scores $0\leq q_{(1)}\leq\cdots\leq q_{(N)}\leq 1$, we define the optimal separation threshold as $j^{\star}\in[N]$ such that
+Let $0=q_{(0)}\leq q_{(1)}\leq\cdots\leq q_{(N)}$, where $\{q_{(i)}\in\mathbb{R}\}_{i=1}^{N}$ is a sorted permutation of $\{\widehat{p}(x_{i})\}_{i=1}^{N}$.
+Further we let $\{x_{(i)}\}_{i=1}^{N}$ and $\{y_{(i)}\}_{i=1}^{N}$ denote the sorted list’s corresponding examples and labels.
+We define the optimal separation threshold as $j^{\star}\in\{0,1,...,N\}$ such that
 $j^{\star}=\operatorname*{arg\,max}_{j}$ (|$\left(\lvert\{y_{(i)}=1:i\leq j\}\rvert-\lvert\{y_{(i)}=2:i\leq j\}\rvert\right)$
-$=\operatorname*{arg\,max}_{j}$ (|$\left(\lvert\{y_{(i)}=2:i>j\}\rvert-\lvert\{y_{(i)}=2:i>j\}\rvert\right).$ (1)
+$=\operatorname*{arg\,max}_{j}$ (|$\left(\lvert\{y_{(i)}=2:i>j\}\rvert-\lvert\{y_{(i)}=1:i>j\}\rvert\right).$ (1)
 In other words, on either side of $j^{\ast}$, it has the largest discrepancy in the number of examples between the two classes.
-This captures the intuition of Figure 1(b) — our goal is to find a threshold that best separates one class from the other.
+This captures the intuition of Figure 2(a) — our goal is to find a threshold that best separates one class from the other.
 We quickly remark that ties are broken by choosing the largest $j^{\star}$ that attains the argmax if class $1$ is the minority class and the lowest $j^{\star}$ otherwise.
@@
 We now provide a reduction of finding $j^{\star}$ to an 1-dimensional agnostic active learning problem.
-We define the hypothesis class =$\mathcal{H}=\{h_{0},h_{1},...,h_{N}\}$ where each hypothesis $h_{j}:[0,1]\rightarrow\{1,2\}$ is defined as
-$h_{j}(q)=\begin{cases}1&\text{if}\,q\leq q_{(j)}\\ 2&\text{if}\,q>q_{(j)}\end{cases}.$
-and $q_{(0)}=0$ defines the hypothesis $h_{0}$ that predicts class $2$ at all times.
+We define the hypothesis class =$\mathcal{H}=\{h_{0},h_{1},...,h_{N}\}$ where each hypothesis $h_{j}$ is defined as $h_{j}(q)=\begin{cases}1&\text{if}\,q\leq q_{(j)}\\ 2&\text{if}\,q>q_{(j)}\end{cases}$.
+Here, $q_{(0)}=0$ defines the hypothesis $h_{0}$ that predicts class $2$ at all times.
 The empirical zero-one loss for each hypothesis is then defined as $\mathcal{L}(h_{j})=\sum_{i=1}^{N}\mathbf{1}\{h_{j}(q_{(i)})\neq y_{(i)}\}$.
 In Appendix A, we show that optimizing for the zero-one loss $\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$ is equivalent to (4.1).
-Namely, $j^{\star}\in\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$, transforming the problem of finding the optimal separation threshold to an 1-D agnostic active learning problem.
+Namely, with ties broken similar to above, $j^{\star}=\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$.
+Figure 3: Visualization of multi-class classification.
+For each class, we formulate the problem as a one-vs-rest binary classification problem by sorting examples based on margin scores.
+The black arrows indicates the optimal separation thresholds for each class.
 Multi-Class Classification.
@@
 (2022).
-Specifically, for each class $k$, one can view the problem of class-$k$ v.s.
+As shown in Figure 3, for each class $k$, we can view the problem of class-$k$ v.s.
 others as a binary classification problem.
-In other words, class $k$ can be seen as class $1$ and classes $\{1,...,k-1,k+1,...,N\}$ can be seen as class $2$.
 The goal therefore becomes finding all $K$ optimal separation thresholds, which is equivalent with solving $K$ 1-D agnostic active learning problems.
-Moreover, let $\widehat{p}:X\rightarrow\Delta^{(K-1)}$ denote the neural network prediction function mapping examples to sigmoid scores.
-Instead of using sigmoid scores, for class $k$, we use the margin scores $q_{i}^{k}:=\max_{k^{\prime}}[\widehat{p}(x_{i})]_{k^{\prime}}-[\widehat{p}(x_{i})]_{k}$ to sort the examples.
-When multiple examples have the same margin score, we further break ties by their corresponding confidence scores $[\widehat{p}(x_{i})]_{k}$.
-We also note that sorting with margin scores is equivalent to sorting with sigmoid scores in the binary classification setting, so this is a strictly more general formulation to multi-class classification.
-  Input: Pool $X$, number of rounds $T$, retraining batch size $B_{\text{train}}$, number of parallel annotations $B_{\text{parallel}}$.
-  Initialize: Uniformly sample $B$ elements without replacement from $X$ to form $L_{0}$. Let $U_{0}\leftarrow X\backslash L_{0}$.
+Moreover, let $\widetilde{p}:X\rightarrow\Delta^{(K-1)}$ denote the neural network prediction function, mapping examples to softmax scores.
+For each class $k$, we use the margin scores $\widehat{p}_{i}^{k}:=[\widetilde{p}(x_{i})]_{k}-\max_{k^{\prime}}[\widetilde{p}(x_{i})]_{k^{\prime}}$ to sort the examples and break ties by their corresponding confidence scores $[\widetilde{p}(x_{i})]_{k}$.
+Formally,
+$\left(q_{(1)}^{k}\leq\mkern-2.0mu{\cdots}\leq q_{(N)}^{k}\mkern-1.0mu{:}\text{ sorted permutation of\,}\{\widehat{p}_{i}^{k}\}_{i=1}^{N}\right)\mkern-1.0mu{\land}\left(q_{(i)}^{k}\mkern-1.0mu{=}q_{(i+1)}^{k}\Rightarrow[\widetilde{p}(x_{i})]_{k}\geq[\widetilde{p}(x_{i+1})]_{k}\right).$ (2)
+Note that sorting by margin scores is equivalent to sorting by sigmoid scores for binary classification.
+  Input: Pool $X$, #Rounds $T$, retraining batch size $B_{\text{train}}$, number of parallel annotations $B_{\text{parallel}}$.
+  Initialize: Uniformly sample $B$ elements from $X$ to form $L_{0}$. Let $U_{0}\leftarrow X\backslash L_{0}$.
   for $t=1,...,T-1$ do
@@
      Find optimal separation thresholds
-     Initialize labeled set $L_{t}\leftarrow L_{t-1}$.
-     Initialize budget per class $b\leftarrow B_{\text{train}}/2K$.
+     Initialize labeled set $L_{t}\leftarrow L_{t-1}$ and budget per class $b\leftarrow B_{\text{train}}/2K$.
      for $k$ in $\text{RandPerm}(\{1,...,K\})$ do
-        Compute margin scores $q_{1}^{k},...,q_{N}^{k}$ based on $f_{t-1}$ and sort them by $q_{(1)}^{k}\leq\cdots\leq q_{(N)}^{k}$ and break ties so that $q_{(i)}^{k}=q_{(i+1)}^{k}\implies[\widehat{p}(x_{(i)})]_{k}\geq[\widehat{p}(x_{(i+1)})]_{k}$.
-        Find optimal separation threshold: $L_{t}\leftarrow\text{VReduce}(L_{t},b,k,B_{\text{parallel}},\{q^{k}_{(i)}\}_{i=1}^{N})$.
+        Sort margin scores $0=q_{(0)}^{k}\leq q_{(1)}^{k}\leq\cdots\leq q_{(N)}^{k}$ based on equation (2).
+        Let $x_{(i)}^{k},y_{(i)}^{k}$ denote the example and label corresponding to $q_{(i)}^{k}$.
+        Identify threshold for class $k$: $L_{t}\leftarrow\text{VReduce}(L_{t},b,k,B_{\text{parallel}},\{(x^{k}_{(i)},y^{k}_{(i)}),q^{k}_{(i)}\}_{i=1}^{N})$.
      end for
-     Annotate rare and uncertain examples
+     Annotate examples around the identified threshold
      Compute budget per class $b\leftarrow(B_{\text{train}}-|L_{t}|)/K$.
      for $k$ in $\text{RandPerm}(\{1,...,K\})$ do
-        Estimate separation threshold
-$\widehat{j}^{k}\leftarrow$ $\operatorname*{arg\,max}_{j}(|\{y_{(i)}=1:x_{(i)}\in L_{t}\,\text{and}\,i\leq j\}|$
-|$-|\{y_{(i)}=1:x_{(i)}\in L_{t}\,\text{and}\,i\leq j\}|)$
-and break ties by choosing the index closest to $\frac{N}{2}$. Annotate $b$ unlabeled examples with sorted indices closest to $\widehat{j}^{k}$ and insert to $L_{t}$.
+        Estimate separation threshold (break ties by choosing the index closest to $\frac{N}{2}$):$\quad\widehat{j}^{k}\leftarrow\operatorname*{arg\,max}_{j}(|\{y_{(i)}=k:x_{(i)}\in L_{t}\,\text{and}\,i\leq j\}|-|\{y_{(i)}\neq k:x_{(i)}\in L_{t}\,\text{and}\,i\leq j\}|)$.
+Annotate $b$ unlabeled examples with sorted indices closest to $\widehat{j}^{k}$ and insert to $L_{t}$.
      end for
@@
 Algorithm 1 DIRECT: DImension REduction for aCTive Learning under Imbalance and Label Noise
-  Input: Labeled set $L$, budget $b$, class of interest $k$, parallel batch size $B_{\text{parallel}}$, examples and ground truth labels sorted by uncertainty $\{q_{(i)}\}_{i=1}^{N}$.
-  Define: For examples in the labeled set, we let $y_{(i)}$ denote the label of the example corresponding to $q_{(i)}$.
-  Initialize: Version space: shortest segment of indices $[I,J]$, with $q_{(I)},q_{(J)}$ $\forall i\leq I,q_{(i)}\in L\implies y_{(i)}=k$ and $\forall j\geq J,q_{(j)}\in L\implies y_{(j)}\neq k$.
+4.2 Algorithm
+We are now ready to state our algorithm DIRECT as shown in Algorithm 1.
+Each round of DIRECT follows a two-phased procedure, where the first phase aims to identify the optimal separation threshold for each class.
+The second phase then annotates examples closest to the estimated optimal separation thresholds for each class.
+We spend half each round’s budget for both phases.
+During the first phase, to identify the optimal separation threshold for all classes, we loop over each class $k$ and run a agnostic active learning procedure for the corresponding 1-D class-k v.s.
+rest reduction.
+Our agnostic active learning subroutine is formally outlined in Algorithm 2, VReduce.
+Our method corresponds to the fixed-budget ACED Katz-Samuels et al.
+(2021) algorithm, adopted to the threshold classifier scenario.
+To explain the intuition about ACED in this scenario, we first note that in the separable case with $\mathcal{L}(h_{j^{\star}})=0$, we could simply run a bisection procedure to learn $j_{k}^{\star}$, the optimal separation threshold for the $k$-th class.
+However, as $j^{\star}$ cannot perfectly separate the classes, ACED circumvents this by maintaining a version space of possible thresholds, namely the interval $[I,J]$.
+Statistically with high likelihood, the optimal separation threshold $j_{k}^{\star}$ lies in this version space of $[I,J]$.
+During each of the $m$ rounds of VReduce, $B_{\text{parallel}}$ samples are annotated and the version space’s length is shrank by a factor of $1/c$.
+The shrinkage rate $c$ is determined by the budget and batch size, so that after the final iteration, the version space has exactly one hypothesis left.
+The second phase of DIRECT simply annotates examples closest to each optimal separation threshold, aiming to annotate a class-balanced and uncertain examples.
+  Input: Labeled set $L$, budget $b$, class of interest $k$, parallel batch size $B_{\text{parallel}}$, examples and ground truth labels with sorted uncertainty $\{x_{(i)}^{k},y_{(i)}^{k},q_{(i)}^{k}\}_{i=1}^{N}$ (unlabeled $y_{(i)}^{k}$ hidden to learner).
+  Initialize: Version space as the shortest segment of indices $[I,J]$ such that: for each labeled example $x_{(i)}$ with $i\leq I$, $y_{(i)}=k$, and for each labeled example $x_{(j)}$ with $j\geq J$, $y_{(j)}\neq k$.
   Initialize: Number of iterations $m\leftarrow\frac{b}{B_{\text{parallel}}}$. Shrinking factor $c\leftarrow\sqrt[m]{J-I}$
   for $t=1,...,m$ do
-     Sample uniformly at random $B_{\text{parallel}}$ unlabeled examples in $q_{(I)}$, …, $q_{(J)}$ for annotation and insert to $L$.
-     Find $[I^{\prime},J^{\prime}]\subset[I,J]$ such that
-$I^{\prime},J^{\prime}=\operatorname*{arg\,min}_{i,j:j-i=\frac{1}{c}(J-I)}\max\{\widehat{\mathcal{L}}^{k}(i),\widehat{\mathcal{L}}^{k}(j)\}\quad\text{where }$
-^$\widehat{\mathcal{L}}^{k}(s)=\sum_{\begin{subarray}{c}r\leq s:\\ q_{(r)}\in L\end{subarray}}\mathbf{1}\{y_{(r)}\neq k\}+\sum_{\begin{subarray}{c}r>s:\\ q_{(r)}\in L\end{subarray}}\mathbf{1}\{y_{(r)}=k\}.$
-     Update version space $I,J\leftarrow I^{\prime},J^{\prime}$.
+     Uniformly sample $B_{\text{parallel}}$ unlabeled examples in $x_{(I)}^{k}$, …, $x_{(J)}^{k}$ for annotation and insert to $L$.
+     For each $0\leq s\leq N$, compute ^$\widehat{\mathcal{L}}^{k}(s)=\sum_{r\leq s:x_{(r)}\in L}\mathbf{1}\{y_{(r)}\neq k\}+\sum_{r>s:x_{(r)}\in L}\mathbf{1}\{y_{(r)}=k\}$.
+     Shrink version space by $\frac{1}{c}$: $I,J\leftarrow\operatorname*{arg\,min}_{i,j\in[I,J]:j-i=\frac{1}{c}(J-I)}\max\{\widehat{\mathcal{L}}^{k}(i),\widehat{\mathcal{L}}^{k}(j)\}$.
   end forReturn: Updated labeled set $L$.
 Algorithm 2 VReduce: Version Space Reduction
-4.2 Algorithm
-We are now ready to state our algorithm as shown in Algorithm 1.
-The algorithm follows a two-phased procedure, where the first phase aims to identify the optimal separation threshold for each class.
-The second phase then annotates examples closest to the estimated optimal separation thresholds for each class.
-Our set of labeled examples at anytime are maintained in $L_{t}$.
-To identify the optimal separation threshold, we loop over each class $k$ and run a fix-budget batch active learning procedure for the one-dimensional reduction (as shown in Algorithm 2) on the set of labeled examples.
-Our method is motivated by the fixed-budget algorithm for active classification, ACED (Katz-Samuels et al., 2021).
-Intuitively, in the non-noisy or separable setting, where the loss of $\mathcal{L}(h_{j^{\star}})=0$, we could run a binary search procedure to learn $j^{\star}$.
-However, we can’t assume that $j^{\star}$ perfectly separates the classes.
-ACED circumvents this by maintaining a version space of threshold classifiers, namely an interval $[I,J]$.
-In each of the $b/B_{\text{parallel}}$ rounds, $B_{\text{parallel}}$ samples are taken and appended to the set of labeled examples $L$.
-Then, an interval $I^{\prime},J^{\prime}$ is computed, whose length is a factor of $1/c$ smaller than the current $I,J$ of low-loss classifiers.
-The shrinkage rate $c$ is determined by the budget and batch size, so that after the final iteration, the version space has exactly one hypothesis left.
-The fixed confidence (or alternatively $\delta$-PAC) analogue of ACED is known to match instance optimal lower bounds for active classification up to logarithmic factors.
-Under benign noise conditions, the probability of misidentifying the optimal separation threshold decays exponentially w.r.t.
-budget $b$.
-To address batch labeling, we note the number of examples collected before retraining is usually far greater than the number of annotators annotating in parallel.
-In particular, we let $B_{\text{train}}$ denote the number of examples the algorithm collects before the neural network is retrained.
+To address batch labeling, we let $B_{\text{train}}$ denote the number of examples the algorithm collects before the neural network is retrained.
 In practice, this number is usually determined by the constraints of computational training cost.
 On the other hand, we let $B_{\text{parallel}}$ denote the number of examples annotated in parallel.
-As discussed above, $B_{\text{parallel}}\ll B_{\text{train}}$ in practice.
-Lastly, as will be discussed in Section 6, a variant of our algorithm can also be proposed for asynchronous labeling.
+We note that, in practice, the number of examples collected before retraining is usually far greater than the number of annotators annotating in parallel, i.e., $B_{\text{parallel}}\ll B_{\text{train}}$.
+Lastly, as will be discussed in Section 6, our algorithm can also be modified for asynchronous labeling.
+Theoretical Comparison with GALAXY.
+As mentioned in Section 3.2, GALAXY’s graph-based approach aims to identify all cuts and sample examples around all of them equally.
+On the other hand, DIRECT aims to identify only the separation threshold and sample around it, which is superior as we have argued before and shown in our results.
+We now present a more theoretical comparison.
+As we show in Appendix A, the graph-based approach in GALAXY will identify and annotate around at least one more cut in addition to the optimal separation threshold, with probability at least $1-\exp(-b\log(\frac{1}{1-\eta})/2)$.
+Here, $b$ is the budget of a single round of annotation and $\eta$ is the label noise ratio (see Appendix A for more details).
+This implies, when the budget $b$ is large, GALAXY will likely annotate around unnecessary cuts.
+This is in contrast with the agnostic active learning approach we take in DIRECT, where as shown by Katz-Samuels et al.
+(2021), the probability of misidentifying the optimal separation threshold decays exponentially w.r.t.
+budget $b$.
+In other words, with a large budget $b$, with high likelihood, DIRECT will focus its annotation around the optimal separation threshold.
+Lastly, time complexity analysis in Appendix C shows DIRECT’s superior speed compared to BADGE and GALAXY.
 5 Experiments
@@
   1. 1.
-    Fine-tuning ResNet-18 with vanilla supervised training on imbalanced datasets similar to Zhang et al. (2022).
+    Supervised fine-tuning of ResNet-18 on imbalanced datasets similar to Zhang et al. (2022).
   2. 2.
-    Fine-tuning large pretrained model (CLIP ViT-B32) with semi-supervised training strategies under the LabelBench framework (Zhang et al., 2023a).
-In Section 5.1, we document these setups in detail.
+    Fine-tuning large pretrained model (CLIP ViT-B32) with semi-supervised training strategies under the LabelBench framework (Zhang et al., 2024a).
 For both evaluation setups, we first evaluate the performance of DIRECT under the noiseless setting in Section 5.1, showing its superior label-efficiency and ability to accommodate batch labeling.
@@
 Under this setting, we also include an ablation study of the performance of DIRECT on various levels of label noises.
-While we highlight many results in this section, see Appendix B for all results.
-5.1 Experiment Setup
-ResNet-18 Experiments.
-ResNet-18 with passive training has been the standard evaluation in existing deep active literature (Ash et al., 2019; Zhang et al., 2022).
-Our experiment setup utilizes the CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), SVHN (Netzer et al., 2011) and PathMNIST (Yang et al., 2021) image classification datasets.
-The original forms of these datasets are roughly balanced across 9, 10 or 100 classes.
-We construct an extremely imbalanced dataset by grouping a large number of classes into one majority class.
-For example, given a balanced dataset above with $M$ classes.
-We generate an imbalanced dataset with $K$ classes ($K<M$) by the first $K-1$ classes from the original dataset and combining the rest of the classes $K,...,M$ into a single majority class $K$.
-Imbalance ratios are shown in Table 1.
-For neural network training, we utilize the standard passive training on labeled examples with cross entropy loss and Adam optimizer (Kingma & Ba, 2014).
-The ResNet-18 model (He et al., 2016) is pretrained on ImageNet (Deng et al., 2009) from the PyTorch library.
-To address data imbalance, we utilize a reweighted cross entropy loss by the inverse frequency of the number of labeled examples in each class.
-For experiments with label noise, we further add a $10\%$ label smoothing during training (Müller et al., 2019).
-LabelBench Experiments.
-Proposed by Zhang et al.
-(2023a), LabelBench evaluates active learning performance in a more comprehensive framework.
-Here, we fine-tune the large pretrained model from CLIP’s ViT-B32 model (Radford et al., 2021).
-The framework also utilizes semi-supervised learning method FlexMatch (Zhang et al., 2021) to further leverage the unlabeled examples in the pool for training.
-We conduct experiments on the two imbalanced datasets in LabelBench, with FMoW (Christie et al., 2018) and iWildcam (Beery et al., 2021).
-Similar to the ResNet-18 experiments, we use a $10\%$ label smoothing in the loss function to improve training under label noise.
-We did find FlexMatch to perform poorly under the combination of imbalance and label noise, so we used the passive training method for label noise experiments.
-Name # Classes $K$ $N$ Imb Ratio $\gamma=\frac{\min_{k}N_{k}}{\max_{k^{\prime}}N_{k^{\prime}}}$
-Imb CIFAR-10 $2$ $50000$$1$.1111.1111.1111
-Imb CIFAR-10 $3$ $50000$1250.1250.1250.1250
-Imb CIFAR-100 $2$ $50000$0101.0101.0101.0101
-Imb CIFAR-100 $3$ $50000$0102.0102.0102.0102
-Imb CIFAR-100 $10$ $50000$0110.0110.0110.0110
-Imb SVHN $2$ $73257$0724.0724.0724.0724
-Imb SVHN $3$ $54448$2546.2546.2546.2546
-PathMNIST $2$ $89996$1166.1166.1166.1166
-FMoW $62$ $76863$0049.0049.0049.0049
-iWildCam $14$ $129809$ $4.57\cdot 10^{-5}$
-Table 1: Dataset settings for our experiments. $N$ denotes the total number of examples in our dataset. $\gamma$ is the class imbalance ratio defined in Section 3.1.
+While we highlight many results in this section, see Appendix E for complete results.
+Experiment Setups.
+Our experiments utilize $10$ imbalanced datasets derived from popular computer vision datasets.
+For the ResNet experiments, we utilize imbalanced versions of CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), SVHN (Netzer et al., 2011) and PathMNIST (Yang et al., 2021) datasets.
+For the LabelBench experiments, we utilize the FMoW (Christie et al., 2018) and iWildcam (Beery et al., 2021) datasets.
+We refer the readers to Appendix D for more details on our experiment setups.
 (a) Imbalanced CIFAR-10, three classes
@@
 (e) FMoW, Balanced Pool Accuracy
-(f) Imbalanced SVHN, two classes, number of labeled examples in minority class
+(f) Imbalanced SVHN, two classes, # of minority labeled examples
 Figure 4: Performance comparison of DIRECT against other baselines algorithms in the noiseless but imbalanced setting.
@@
 $B_{\text{parallel}}=1$ is equivalent with sychornous labeling.
-Results are averaged over four trials and the shaded areas represent standard errors around the average performances.
+Results are averaged over four trials and the shaded areas represent standard errors around the mean.
 5.1 Experiments under Imbalance, without Label Noise
@@
 The rest of the algorithms primarily optimizes for different notions of informativeness such as diversity and uncertainty.
-For the LabelBench experiments, due to the large dataset and model embedding sizes, we choose a subset of the algorithms that are computationally efficient and among top performers in the ResNet-18 results.
-In particular, we compared against BADGE, Margin Sampling (Balcan et al., 2006; Kremer et al., 2014), CORESET and GALAXY.
-As highlighted in Figures 1(a) and 3(a)-(d), DIRECT consistently and significantly outperforms existing algorithms on the ResNet-18 experiments.
-In Figures 1(c) and 3(e), we demonstrate the increased label-efficiency is also consistently shown in the LabelBench experiments.
+For the LabelBench experiments, due to the large dataset and model embedding sizes, we choose a subset of the algorithms that are computationally efficient and among top performers in the ResNet-18 results, including BADGE, Margin Sampling, CORESET and GALAXY.
+As highlighted in Figures 1(a) and 4(a)-(d), DIRECT consistently and significantly outperforms existing algorithms on the ResNet-18 experiments.
+In Figures 1(c) and 4(e), we demonstrate the increased label-efficiency is also consistently shown in the LabelBench experiments.
 Compared to random sampling, DIRECT can save more than 80% of the annotation cost on imbalanced SVHN experiment of Figure 4(c).
@@
 This is in comparison to the synchoronous nature of GALAXY, where it is always using $B_{\text{parallel}}=1$.
-On Figures 1(a) and 4(a), we see that DIRECT outperforms GALAXY with synchronous labeling.
+On Figures 1(a) and 5(a), we see that DIRECT outperforms GALAXY with synchronous labeling.
 Furthermore, in these experiments we also see using $B_{\text{parallel}}=5$ only affects algorithm performances minimally for DIRECT.
@@
 (a)-(b) are balanced accuracy on ResNet-18 experiments while (c) shows results under the LabelBench framework.
-Results are averaged over four trials and the shaded areas represent standard errors around the average performances.
+Results are averaged over four trials and the shaded areas represent standard errors around the mean.
 (a) Imbalanced CIFAR-100, two classes, no label noise
@@
 (c) Imbalanced CIFAR-100, two classes, 15% label noise
-Figure 5: Performance of DIRECT against baseline algorithms under different levels of label noise.
-Results are averaged over four trials and the shaded areas represent standard errors around the average performances.
+Figure 6: Performance of DIRECT against baseline algorithms under different levels of label noise. Results are averaged over four trials and the shaded areas represent standard errors around the mean.
 5.2 Experiments under Imbalance and Label Noise
@@
 In fact, the noiseless experiment in Figure 6(a) is the only setting DIRECT slightly underperforms GALAXY in terms of generalization accuracy.
-However, we see DIRECT becomes more label-efficient under the more practical label noise settings.
+However, we see DIRECT becomes more label-efficient under label noise.
 It is also worth noting that with high label noise of $20\%$, we observe in Figure 1(b) that existing algorithms underperform random sampling.
@@
 In this work, we also addressed the batch sampling problem of current SOTA algorithm, GALAXY (Zhang et al., 2022), by annotating multiple examples in parallel.
-Studying asynchronous labeling could a natural extension of our work.
-We think a potential solution is to utilize an asynchronous variant of one-dimensional active learning algorithm.
+Studying asynchronous labeling could be a natural extension of our work.
+A potential solution is to utilize an asynchronous variant of one-dimensional active learning algorithm.
 In addition, one can further batch the labeling process across different classes to further accommodate an even larger number of parallel annotators.
+Lastly, we refer the readers to Appendix F for our impact statement.
 Acknowledgements
 This work has been supported in part by NSF Award 2112471.
-Impact Statement
-In the rapidly evolving landscape of machine learning, the efficacy of active learning in addressing data imbalance and label noise is a significant stride towards more robust and equitable AI systems.
-This research explores how active learning can effectively mitigate the challenges posed by imbalanced datasets and erroneous labels, prevalent in real-world scenarios.
-The positive impacts of this research are multifaceted.
-It enhances the accessibility and utility of machine learning in domains where data imbalance is a common challenge, such as healthcare, finance, and social media analytics.
-By improving class-balancedness in annotated sets, models trained on these datasets are less biased and more representative of real-world distributions, leading to fairer and more accurate outcomes.
-Additionally, this research contributes to reducing the time and cost associated with data annotation, which is particularly beneficial in fields where expert annotation is expensive or scarce.
-However, if not carefully implemented, active learning strategies could inadvertently introduce new biases or amplify existing ones, particularly in scenarios where the initial data is severely imbalanced or contains deeply ingrained biases.
-Furthermore, the advanced nature of these techniques may widen the gap between organizations with access to state-of-the-art technology and those without, potentially exacerbating existing inequalities in technology deployment.
 Appendix A Equivalent Objective
@@
 $=\operatorname*{arg\,max}_{j}\sum_{i=1}^{j}\left(\mathbf{1}\{y_{(i)}=1\}-\mathbf{1}\{y_{(i)}=2\}\right)$
-Appendix B All Results
-B.1 Noiseless Results under Imbalance
+Appendix B Theoretical Analysis
+In this section, we analyze the performance of GALAXY under random label noise and show the probability of identifying and sampling around additional cuts increases as more examples are labeled.
+This is in contrast to the DIRECT’s agnostic active learning approach, where the probability of identifying and sampling around only the optimal separation threshold decays exponentially in the number of labeling budget.
+Specifically, under the binary classification scenario, one is given a sorted list of $N$ examples $\{x_{(i)}\}_{i=1}^{N}$, with ground truth labels $y_{(1)}^{\star}=y_{(2)}^{\star}=...=y_{(N_{1})}^{\star}=1$ and $y_{(N_{1}+1)}^{\star}=...=y_{(N_{1}+N_{2})}^{\star}=2$, where $N_{1}+N_{2}=N$.
+Under uniform i.i.d.
+label noise with noise ratio $\eta>0$, the observed labels are denoted as $\{y_{(i)}\}_{i=1}^{N}$, where $\mathbb{P}(y_{(i)}\neq y_{(i)}^{\star})=\eta$.
+In other words, the observed label is flipped with probability $\eta$.
+Theorem B.1.
+Given a budget of $b>2\log N$, let $M_{b}$ be the random variable denoting number of identified cuts in addition to the optimal separation threshold by one round of GALAXY.
+We must have $\mathbb{P}(M_{b}\geq 1)\geq 1-\exp(-b\log(\frac{1}{1-\eta})/2)$, implying GALAXY samples around at least one more cut in addition to the optimal separation threshold with high probability.
+Proof.
+In the perfect scenario where GALAXY does not receive any corrupted labels, it would use $\log N$ budget with bisection to find the optimal separation threshold and annotate around it.
+However, within the first $\frac{b}{2}$ annotations, whenever GALAXY receives a corrupted label, it will identify a cut in addition to the optimal separation threshold, i.e., $M_{b}\geq 1$.
+Therefore, the probability of $M_{b}\geq 1$ is greater than the probability of receiving at least one corrupted labels in the first $\frac{b}{2}$ annotations.
+With simple probability bound, we can show that
+$\mathbb{P}(M_{b}\geq 1)>1-(1-\eta)^{b/2}=1-\exp(b\log(1-\eta)/2)=1-\exp(-b\log(\frac{1}{1-\eta})/2).$
+As the theorem suggests, when $b$ is large, GALAXY will identify and annotate around at least one additional cut with high probability.
+Appendix C Time Complexity
+The computation complexity for each batch of DIRECT is $O(KN\log(N)+B_{\text{train}}N)$ for data selection plus the training and inference costs of the neural network.
+$O(KN\log(N))$ comes from sorting examples by their margin scores for each class and $O(B_{\text{train}}N)$ is the cost for running Algorithm 2 for $O(B_{\text{train}})$ iterations.
+Each iteration of Algorithm 2 only costs $O(N)$ time as we can efficiently solve the objective by cumulative sums.
+We note that the cost associated with neural network training and inference is always the dominating factor.
+For comparisons, BADGE has time complexity $O(B_{\text{train}}N(K+D))$, significantly more expensive than DIRECT, with $D$ denotes the dimensionality of the penultimate layer features.
+In addition, GALAXY has computational complexity of $O(KN\log(N))+B_{\text{train}}KN$, also more expensive than DIRECT.
+In all of our experiments, both BADGE and GALAXY indeed is slower than DIRECT.
+We further note that the time complexity factor of $K$ in DIRECT can be easily parallelized by conducting the $K$ sorting procedures on different CPU cores.
+In total, our experiments are conducted on NVIDIA 3090 ti GPUs. Each trial of the ResNet-18 experiment takes less than two hours while each trial of the LabelBench experiments takes roughly 12 hours.
+Appendix D Experiment Setup
+Name $K$ $N$ Imb Ratio $\gamma=\frac{\min_{k}N_{k}}{\max_{k^{\prime}}N_{k^{\prime}}}$
+Imb CIFAR-10 $2$ $50000$$1$.1111.1111.1111
+Imb CIFAR-10 $3$ $50000$1250.1250.1250.1250
+Imb CIFAR-100 $2$ $50000$0101.0101.0101.0101
+Imb CIFAR-100 $3$ $50000$0102.0102.0102.0102
+Imb CIFAR-100 $10$ $50000$0110.0110.0110.0110
+Imb SVHN $2$ $73257$0724.0724.0724.0724
+Imb SVHN $3$ $54448$2546.2546.2546.2546
+PathMNIST $2$ $89996$1166.1166.1166.1166
+FMoW $62$ $76863$0049.0049.0049.0049
+iWildCam $14$ $129809$ $4.57\cdot 10^{-5}$
+Table 1: Dataset settings for our experiments. $N$ denotes the total number of examples in our dataset. $\gamma$ is the class imbalance ratio defined in Section 3.1.
+ResNet-18 Experiments.
+ResNet-18 with passive training has been the standard evaluation in existing deep active literature [Ash et al., 2019, Zhang et al., 2022].
+Our experiment setup utilizes the CIFAR-10, CIFAR-100 [Krizhevsky et al., 2009], SVHN [Netzer et al., 2011] and PathMNIST [Yang et al., 2021] image classification datasets.
+The original forms of these datasets are roughly balanced across 9, 10 or 100 classes.
+We construct an extremely imbalanced dataset by grouping a large number of classes into one majority class.
+For example, given a balanced dataset above with $M$ classes.
+We generate an imbalanced dataset with $K$ classes ($K<M$) by the first $K-1$ classes from the original dataset and combining the rest of the classes $K,...,M$ into a single majority class $K$.
+Imbalance ratios are shown in Table 1.
+For neural network training, we utilize the standard passive training on labeled examples with cross entropy loss and Adam optimizer [Kingma and Ba, 2014].
+The ResNet-18 model [He et al., 2016] is pretrained on ImageNet [Deng et al., 2009] from the PyTorch library.
+To address data imbalance, for all algorithms, we utilize a reweighted cross entropy loss by the inverse frequency of the number of labeled examples in each class.
+For experiments with label noise, we further add a $10\%$ label smoothing during training [Müller et al., 2019] for all algorithms.
+LabelBench Experiments.
+Proposed by Zhang et al.
+[2024a], LabelBench evaluates active learning performance in a more comprehensive framework.
+Here, we fine-tune the large pretrained model from CLIP’s ViT-B32 model [Radford et al., 2021].
+The framework also utilizes semi-supervised learning method FlexMatch [Zhang et al., 2021] to further leverage the unlabeled examples in the pool for training.
+We conduct experiments on the two imbalanced datasets in LabelBench, with FMoW [Christie et al., 2018] and iWildcam [Beery et al., 2021].
+Similar to the ResNet-18 experiments, for all algorithms, we use a $10\%$ label smoothing in the loss function to improve training under label noise.
+We did find FlexMatch to perform poorly under the combination of imbalance and label noise, so we used the passive training method for label noise experiments.
+Appendix E All Results
+E.1 Noiseless Results under Imbalance
 Figure 7: Imbalanced CIFAR-10, two classes.
@@
 Figure 15: LabelBench results in the noiseless setting.
-B.2 Label Noise Results under Imbalance
+E.2 Label Noise Results under Imbalance
 Figure 16: Imbalanced CIFAR-10, two classes, 10% label noise.
@@
 Figure 24: LabelBench results in the 10% label noise setting.
+Appendix F Impact Statement
+In the rapidly evolving landscape of machine learning, the efficacy of active learning in addressing data imbalance and label noise is a significant stride towards more robust and equitable AI systems.
+This research explores how active learning can effectively mitigate the challenges posed by imbalanced datasets and erroneous labels, prevalent in real-world scenarios.
+The positive impacts of this research are multifaceted.
+It enhances the accessibility and utility of machine learning in domains where data imbalance is a common challenge, such as healthcare, finance, and social media analytics.
+By improving class-balancedness in annotated sets, models trained on these datasets are less biased and more representative of real-world distributions, leading to fairer and more accurate outcomes.
+Additionally, this research contributes to reducing the time and cost associated with data annotation, which is particularly beneficial in fields where expert annotation is expensive or scarce.
+However, if not carefully implemented, active learning strategies could inadvertently introduce new biases or amplify existing ones, particularly in scenarios where the initial data is severely imbalanced or contains deeply ingrained biases.
+Furthermore, the advanced nature of these techniques may widen the gap between organizations with access to state-of-the-art technology and those without, potentially exacerbating existing inequalities in technology deployment.

### v4

@@
-DIRECT: Deep Active Learning under Imbalance and Label Noise
+Improved Algorithm for Deep Active Learning under Imbalance via Optimal Separation
 Abstract
-Class imbalance is a prevalent issue in real world machine learning applications, often leading to poor performance in rare and minority classes.
-With an abundance of wild unlabeled data, active learning is perhaps the most effective technique in solving the problem at its root – collecting a more balanced and informative set of labeled examples during annotation.
-Label noise is another common issue in data annotation jobs, which is especially challenging for active learning methods.
-In this work, we conduct the first study of active learning under both class imbalance and label noise.
-We propose a novel algorithm that robustly identifies the class separation threshold and annotates the most uncertain examples that are closest from it.
-Through a novel reduction to one-dimensional active learning, our algorithm DIRECT is able to leverage classic active learning theory and methods to address issues such as batch labeling and tolerance towards label noise.
-We present extensive experiments on imbalanced datasets with and without label noise.
-Our results demonstrate that DIRECT can save more than 60% of the annotation budget compared to state-of-art active learning algorithms and more than 80% of annotation budget compared to random sampling.
+Class imbalance severely impacts machine learning performance on minority classes in real-world applications.
+While various solutions exist, active learning offers a fundamental fix by strategically collecting balanced, informative labeled examples from abundant unlabeled data.
+We introduce DIRECT, an algorithm that identifies class separation boundaries and selects the most uncertain nearby examples for annotation.
+By reducing the problem to one-dimensional active learning, DIRECT leverages established theory to handle batch labeling and label noise – another common challenge in data annotation that particularly affects active learning methods.
+Our work presents the first comprehensive study of active learning under both class imbalance and label noise.
+Extensive experiments on imbalanced datasets show DIRECT reduces annotation costs by over 60% compared to state-of-the-art active learning methods and over 80% versus random sampling, while maintaining robustness to label noise.
+Deep Learning, Active Labeling, Class Imbalance
 1 Introduction
@@
 To the best of our knowledge, this is the first deep active learning study to address the challenging yet prevalent scenario where both imbalance and label noise coexist.
+(a) Imbalanced CIFAR-10, two classes, no label noise.
+(b) Imbalanced CIFAR-100, two classes, 20% label noise.
+(c) LabelBench FMoW (62 classes with imbalance), no label noise
+Figure 1: Performance of DIRECT over existing baselines for both noiseless and noisy settings.
+The x-axis represents the total number of labeled examples so far and the y-axis shows the neural network’s balanced accuracy.
+Both (a) and (b) are using supervised training of ResNet-18.
+In (c), we finetune CLIP ViT-B32 model in combination of semi-supervised training under the LabelBench framework (Zhang et al., 2024a).
+$B_{\text{parallel}}$ is the batch size indicating the number of parallel annotators.
+$B_{\text{parallel}}=1$ indicates the synchronous annotation requirement by GALAXY.
+Our algorithm DIRECT takes pre-specified $B_{\text{parallel}}$ as input, which is determined by real world scenarios.
 To bridge the gap between the imbalanced deep active learning and the agnostic active learning literature, we propose a novel reduction of the imbalanced classification problem into a set of one-dimensional agnostic active learning problems.
@@
 On imbalanced datasets, DIRECT achieves state-of-art label efficiency on both supervised fine-tuning of ResNet-18 and semi-supervised fine-tuning of large pretrained model under the LabelBench (Zhang et al., 2024a) framework.
-(a) Imbalanced CIFAR-10, two classes, no label noise.
-(b) Imbalanced CIFAR-100, two classes, 20% label noise.
-(c) LabelBench FMoW (62 classes with imbalance), no label noise
-Figure 1: Performance of DIRECT over existing baselines for both noiseless and noisy settings.
-The x-axis represents the total number of labeled examples so far and the y-axis shows the neural network’s balanced accuracy.
-Both (a) and (b) are using supervised training of ResNet-18.
-In (c), we finetune CLIP ViT-B32 model in combination of semi-supervised training under the LabelBench framework (Zhang et al., 2024a).
-$B_{\text{parallel}}$ is the batch size indicating the number of parallel annotators.
-$B_{\text{parallel}}=1$ indicates the synchronous annotation requirement by GALAXY.
-Our algorithm DIRECT takes pre-specified $B_{\text{parallel}}$ as input, which is determined by real world scenarios.
 To summarize our main contributions:
@@
   • Compared to state-of-art algorithm GALAXY (Zhang et al., 2022), DIRECT allows parallel annotation by multiple annotators while still maintaining significant label-efficiency improvement.
-  • We conduct experiments across eight dataset settings, four levels of label noise and for both ResNet-18 and large pretrained model (CLIP ViT-B32).
+  • We conduct experiments across 12 dataset settings, four levels of label noise and for both ResNet-18 and large pretrained model (CLIP ViT-B32).
 DIRECT consistently outperforms existing baseline algorithms by saving more than $60\%$ annotation cost compared to the best existing algorithm, and more than $80\%$ annotation cost compared to random sampling.
 2 Related Work
-Class-Balanced Deep Active Learning Active earning strategies sequentially and adaptively choose examples for annotation.
-Many uncertainty-based deep active learning methods extend the traditional active learning literature such as margin, least confidence and entropy sampling (Tong and Koller, 2001; Settles, 2009; Balcan et al., 2006; Kremer et al., 2014).
+Class-Balanced Deep Active Learning Active learning strategies sequentially and adaptively choose examples for annotation.
+Many uncertainty-based deep active learning methods extend the traditional active learning literature such as margin, least confidence and entropy sampling (Tong & Koller, 2001; Settles, 2009; Balcan et al., 2006; Kremer et al., 2014).
 These methods have been shown to perform among the top when fine-tuning large pretrained models and combined with semi-supervised learning algorithms (Zhang et al., 2024a).
-More sophisticated methods have been proposed to optimize chosen examples’ uncertainty (Gal et al., 2017; Ducoffe and Precioso, 2018; Beluch et al., 2018), diversity (Sener and Savarese, 2017; Geifman and El-Yaniv, 2017; Citovsky et al., 2021), or a mix of both (Ash et al., 2019, 2021; Wang et al., 2021; Elenter et al., 2022; Mohamadi et al., 2022).
-However, these methods often perform poorly under prevalent and realistic scenarios such as label noises (Khosla et al., 2022) or class imbalance Kothawade et al.
-(2021); Zhang et al.
-(2022, 2024a).
+More sophisticated methods have been proposed to optimize chosen examples’ uncertainty (Gal et al., 2017; Ducoffe & Precioso, 2018; Beluch et al., 2018), diversity (Sener & Savarese, 2017; Geifman & El-Yaniv, 2017; Citovsky et al., 2021), or a mix of both (Ash et al., 2019, 2021; Wang et al., 2021; Elenter et al., 2022; Mohamadi et al., 2022).
+However, these methods often perform poorly under prevalent and realistic scenarios such as label noises (Khosla et al., 2022) or class imbalance (Kothawade et al., 2021; Zhang et al., 2022, 2024a).
 Deep Active Learning under Imbalance Data imbalance and rare instances are prevalent in almost all modern machine learning applications.
-Active learning techniques are effective in addressing the problem in its root by collecting a more class-balanced dataset (Aggarwal et al., 2020; Kothawade et al., 2021; Emam et al., 2021; Zhang et al., 2022; Coleman et al., 2022; Jin et al., 2022; Cai, 2022; Zhang et al., 2024b).
+Active learning techniques are effective in addressing the problem in its root by collecting a more class-balanced dataset (Aggarwal et al., 2020; Kothawade et al., 2021; Emam et al., 2021; Zhang et al., 2022; Coleman et al., 2022; Jin et al., 2022; Cai, 2022; Zhang et al., 2024b; Xie et al., 2024).
 To this end, Kothawade et al.
@@
 Several approaches have been proposed for computing such sampling distributions, e.g.
-Jain and Jamieson (2019); Katz-Samuels et al.
+Jain & Jamieson (2019); Katz-Samuels et al.
 (2020, 2021); Huang et al.
@@
 This is fundamentally different from our work, where our annotator may provide noisy labels.
-Another line of more theoretical active learning research studies active learning with multiple annotators with different qualities (Zhang and Chaudhuri, 2015; Chen et al., 2022).
+Another line of more theoretical active learning research studies active learning with multiple annotators with different qualities (Zhang & Chaudhuri, 2015; Chen et al., 2022).
 The primary goal in these work is to identify examples a weak annotator and a strong annotator may disagree, in order to only use the strong annotator on such instances.
@@
 We quickly remark that ties are broken by choosing the largest $j^{\star}$ that attains the argmax if class $1$ is the minority class and the lowest $j^{\star}$ otherwise.
+Figure 3: Visualization of multi-class classification.
+For each class, we formulate the problem as a one-vs-rest binary classification problem by sorting examples based on margin scores.
+The black arrows indicates the optimal separation thresholds for each class.
 1D Reduction.
@@
 The empirical zero-one loss for each hypothesis is then defined as $\mathcal{L}(h_{j})=\sum_{i=1}^{N}\mathbf{1}\{h_{j}(q_{(i)})\neq y_{(i)}\}$.
-In Appendix A, we show that optimizing for the zero-one loss $\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$ is equivalent to (4.1).
+In Appendix A, we show that optimizing for the zero-one loss $\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$ is equivalent to equation 4.1.
 Namely, with ties broken similar to above, $j^{\star}=\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$.
-Figure 3: Visualization of multi-class classification.
-For each class, we formulate the problem as a one-vs-rest binary classification problem by sorting examples based on margin scores.
-The black arrows indicates the optimal separation thresholds for each class.
 Multi-Class Classification.
@@
 Formally,
-$\left(q_{(1)}^{k}\leq\mkern-2.0mu{\cdots}\leq q_{(N)}^{k}\mkern-1.0mu{:}\text{ sorted permutation of\,}\{\widehat{p}_{i}^{k}\}_{i=1}^{N}\right)\mkern-1.0mu{\land}\left(q_{(i)}^{k}\mkern-1.0mu{=}q_{(i+1)}^{k}\Rightarrow[\widetilde{p}(x_{i})]_{k}\geq[\widetilde{p}(x_{i+1})]_{k}\right).$ (2)
+$\left(q_{(1)}^{k}\leq\mkern-2.0mu{\cdots}\leq q_{(N)}^{k}\mkern-1.0mu{:}\text{ sorted permutation of\,}\{\widehat{p}_{i}^{k}\}_{i=1}^{N}\right)\mkern-1.0mu{\land}$
+$\left(q_{(i)}^{k}\mkern-1.0mu{=}q_{(i+1)}^{k}\Rightarrow[\widetilde{p}(x_{i})]_{k}\geq[\widetilde{p}(x_{i+1})]_{k}\right).$ (2)
 Note that sorting by margin scores is equivalent to sorting by sigmoid scores for binary classification.
+4.2 One-Dimensional Agnostic Active Learning
+The key insight of our approach is to leverage the well-established theory of agnostic active learning for threshold classifiers, which provides robust guarantees even under label noise.
+This allows us to robustly identify the optimal separation threshold $j^{\star}$ from the 1-D reduction above.
+Problem Formulation and Noise Handling.
+In the agnostic setting, we have a sorted sequence $q_{(1)}\leq q_{(2)}\leq\cdots\leq q_{(N)}$ with corresponding (possibly noisy) binary labels $y_{(1)},y_{(2)},\ldots,y_{(N)}$.
+The fundamental challenge is that no hypothesis $h_{j}\in\mathcal{H}$ may achieve zero empirical loss $\mathcal{L}(h_{j})$ due to label noise or model misspecification.
+To formalize how our algorithm handles label noise, we consider the underlying conditional probability function $P(y_{i}|x_{i})$ for each data example $x_{i}$.
+Through the dimensionality reduction $x_{i}\rightarrow q_{i}$ (where $q_{i}$ is the real-valued sigmoid/margin score), we obtain an ordered set of 1-dimensional features $\{q_{i}\}_{i=1}^{N}$.
+This mapping induces a distribution $P(y_{i}|q_{i})$ over the reduced space, which naturally encodes any label noise present in the annotations.
+Our objective is to find the threshold classifier $h_{j^{\star}}$ from the set of 1-dimensional threshold classifiers $\{h_{j}\}$ that minimizes the probability of error with respect to $P(y_{i}|q_{i})$.
+Crucially, our agnostic approach makes no assumptions about the form of $P(y_{i}|x_{i})$ or the induced $P(y_{i}|q_{i})$, allowing the algorithm to handle arbitrary noise models without requiring prior knowledge of the noise distribution.
+Version Space Reduction.
+The VReduce algorithm (Algorithm 1) implements a version space approach that maintains an interval $[I,J]$ representing plausible optimal thresholds.
+The key principle is that if we observe labeled examples $x_{(i)}$ with $i\leq I$ mostly belong to class $1$, and examples $x_{(j)}$ with $j\geq J$ mostly belong to class $2$, then the optimal threshold $j^{\star}$ with high likelihood will lie within $[I,J]$.
+The algorithm proceeds iteratively by: (1) maintaining and updating the version space interval $[I,J]$ based on observed labels, (2) sampling unlabeled examples uniformly within this interval to maximize information gain, (3) shrinking the version space based on empirical loss estimates that account for the noisy observations, and (4) repeating until the labeling budget is exhausted.
+  Input: Labeled set $L$, budget $b$, class of interest $k$, parallel batch size $B_{\text{parallel}}$, sorted examples $\{x_{(i)}^{k},y_{(i)}^{k},q_{(i)}^{k}\}_{i=1}^{N}$ (note $y_{(i)}^{k}$ of unlabeled examples are hidden to the learner).
+  Initialize: Version space $[I,J]$ as the shortest interval such that: $\forall i\leq I$ with $x_{(i)}\in L$: $y_{(i)}=k$, and $\forall j\geq J$ with $x_{(j)}\in L$: $y_{(j)}\neq k$.
+  Number of iterations $m\leftarrow b/B_{\text{parallel}}$. Shrinking factor $c\leftarrow(J-I)^{1/m}$.
+  for $t=1,...,m$ do
+     Sample $B_{\text{parallel}}$ unlabeled examples uniformly from $\{x_{(I)}^{k},\ldots,x_{(J)}^{k}\}$ and add to $L$.
+     Compute empirical loss: ^$\widehat{\mathcal{L}}^{k}(s)=\sum_{r\leq s:x_{(r)}\in L}\mathbf{1}\{y_{(r)}\neq k\}+\sum_{r>s:x_{(r)}\in L}\mathbf{1}\{y_{(r)}=k\}$.
+     Update version space:$[I,J]\leftarrow\operatorname*{arg\,min}_{[i,j]:j-i=(J-I)/c}\max\{\widehat{\mathcal{L}}^{k}(i),\widehat{\mathcal{L}}^{k}(j)\}$.
+  end forReturn: Updated labeled set $L$.
+Algorithm 1 VReduce: Version Space Reduction for Threshold Learning
+Theoretical Guarantees.
+The VReduce algorithm inherits robust theoretical properties from the agnostic active learning literature, achieving near-minimax and instance-optimal sample complexity bounds (Dasgupta et al., 2007; Hanneke et al., 2014).
+Unlike bisection-based methods that fail under label corruption, the disagreement-based framework (Balcan et al., 2006) provides natural robustness with high probability guarantees regardless of the underlying data distribution or noise model.
+Building on the ACED framework (Katz-Samuels et al., 2021), our algorithm extends these classical guarantees to practical batch settings through the $B_{\text{parallel}}$ parameter while preserving all theoretical properties, making it suitable for real-world annotation scenarios with multiple parallel annotators.
+4.3 Algorithm
+We are now ready to state our algorithm DIRECT as shown in Algorithm 2.
+Each round of DIRECT follows a two-phased procedure, where the first phase aims to identify the optimal separation threshold for each class using the agnostic active learning approach described above.
+The second phase then annotates examples closest to the estimated optimal separation thresholds for each class.
+We spend half each round’s budget for both phases.
   Input: Pool $X$, #Rounds $T$, retraining batch size $B_{\text{train}}$, number of parallel annotations $B_{\text{parallel}}$.
@@
      for $k$ in $\text{RandPerm}(\{1,...,K\})$ do
-        Sort margin scores $0=q_{(0)}^{k}\leq q_{(1)}^{k}\leq\cdots\leq q_{(N)}^{k}$ based on equation (2).
+        Sort margin scores $0=q_{(0)}^{k}\leq q_{(1)}^{k}\leq\cdots\leq q_{(N)}^{k}$ based on equation 4.1.
         Let $x_{(i)}^{k},y_{(i)}^{k}$ denote the example and label corresponding to $q_{(i)}^{k}$.
-        Identify threshold for class $k$: $L_{t}\leftarrow\text{VReduce}(L_{t},b,k,B_{\text{parallel}},\{(x^{k}_{(i)},y^{k}_{(i)}),q^{k}_{(i)}\}_{i=1}^{N})$.
+        Identify threshold for class $k$: $L_{t}\leftarrow\text{VReduce}(L_{t},b,k,B_{\text{parallel}},\{x^{k}_{(i)},y^{k}_{(i)},q^{k}_{(i)}\}_{i=1}^{N})$.
      end for
@@
 Algorithm 2 DIRECT: DImension REduction for aCTive Learning under Imbalance and Label Noise
-4.2 Algorithm
-We are now ready to state our algorithm DIRECT as shown in Algorithm 1.
-Each round of DIRECT follows a two-phased procedure, where the first phase aims to identify the optimal separation threshold for each class.
-The second phase then annotates examples closest to the estimated optimal separation thresholds for each class.
-We spend half each round’s budget for both phases.
-During the first phase, to identify the optimal separation threshold for all classes, we loop over each class $k$ and run a agnostic active learning procedure for the corresponding 1-D class-k v.s.
+(a) Imbalanced CIFAR-10, three classes
+(b) Imbalanced CIFAR-100, ten classes
+(c) Imbalanced SVHN, two classes
+(d) PathMNIST, two classes
+(e) FMoW, Balanced Pool Accuracy
+(f) Imbalanced SVHN, two classes, # of minority labeled examples
+Figure 4: Performance comparison of DIRECT against other baselines algorithms in the noiseless but imbalanced setting.
+(a)-(d) are balanced accuracy on ResNet-18 experiments while (e) shows experiment under the LabelBench framework.
+$B_{\text{parallel}}$ indicates the number of parallel annotations as mentioned in Section 4.3.
+$B_{\text{parallel}}=1$ is equivalent with sychornous labeling.
+Results are averaged over four trials and the shaded areas represent standard errors around the mean.
+During the first phase, to identify the optimal separation threshold for all classes, we loop over each class $k$ and run the agnostic active learning procedure VReduce for the corresponding 1-D class-k v.s.
 rest reduction.
-Our agnostic active learning subroutine is formally outlined in Algorithm 2, VReduce.
-Our method corresponds to the fixed-budget ACED Katz-Samuels et al.
-(2021) algorithm, adopted to the threshold classifier scenario.
-To explain the intuition about ACED in this scenario, we first note that in the separable case with $\mathcal{L}(h_{j^{\star}})=0$, we could simply run a bisection procedure to learn $j_{k}^{\star}$, the optimal separation threshold for the $k$-th class.
-However, as $j^{\star}$ cannot perfectly separate the classes, ACED circumvents this by maintaining a version space of possible thresholds, namely the interval $[I,J]$.
-Statistically with high likelihood, the optimal separation threshold $j_{k}^{\star}$ lies in this version space of $[I,J]$.
-During each of the $m$ rounds of VReduce, $B_{\text{parallel}}$ samples are annotated and the version space’s length is shrank by a factor of $1/c$.
-The shrinkage rate $c$ is determined by the budget and batch size, so that after the final iteration, the version space has exactly one hypothesis left.
 The second phase of DIRECT simply annotates examples closest to each optimal separation threshold, aiming to annotate a class-balanced and uncertain examples.
-  Input: Labeled set $L$, budget $b$, class of interest $k$, parallel batch size $B_{\text{parallel}}$, examples and ground truth labels with sorted uncertainty $\{x_{(i)}^{k},y_{(i)}^{k},q_{(i)}^{k}\}_{i=1}^{N}$ (unlabeled $y_{(i)}^{k}$ hidden to learner).
-  Initialize: Version space as the shortest segment of indices $[I,J]$ such that: for each labeled example $x_{(i)}$ with $i\leq I$, $y_{(i)}=k$, and for each labeled example $x_{(j)}$ with $j\geq J$, $y_{(j)}\neq k$.
-  Initialize: Number of iterations $m\leftarrow\frac{b}{B_{\text{parallel}}}$. Shrinking factor $c\leftarrow\sqrt[m]{J-I}$
-  for $t=1,...,m$ do
-     Uniformly sample $B_{\text{parallel}}$ unlabeled examples in $x_{(I)}^{k}$, …, $x_{(J)}^{k}$ for annotation and insert to $L$.
-     For each $0\leq s\leq N$, compute ^$\widehat{\mathcal{L}}^{k}(s)=\sum_{r\leq s:x_{(r)}\in L}\mathbf{1}\{y_{(r)}\neq k\}+\sum_{r>s:x_{(r)}\in L}\mathbf{1}\{y_{(r)}=k\}$.
-     Shrink version space by $\frac{1}{c}$: $I,J\leftarrow\operatorname*{arg\,min}_{i,j\in[I,J]:j-i=\frac{1}{c}(J-I)}\max\{\widehat{\mathcal{L}}^{k}(i),\widehat{\mathcal{L}}^{k}(j)\}$.
-  end forReturn: Updated labeled set $L$.
-Algorithm 2 VReduce: Version Space Reduction
 To address batch labeling, we let $B_{\text{train}}$ denote the number of examples the algorithm collects before the neural network is retrained.
@@
 While we highlight many results in this section, see Appendix E for complete results.
+Our implementation is publicly available at: https://github.com/EfficientTraining/LabelBench/blob/main/LabelBench/strategy/strategy_impl/direct.py
 Experiment Setups.
-Our experiments utilize $10$ imbalanced datasets derived from popular computer vision datasets.
-For the ResNet experiments, we utilize imbalanced versions of CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), SVHN (Netzer et al., 2011) and PathMNIST (Yang et al., 2021) datasets.
-For the LabelBench experiments, we utilize the FMoW (Christie et al., 2018) and iWildcam (Beery et al., 2021) datasets.
+Our experiments utilize $14$ imbalanced datasets derived from popular computer vision datasets.
+For the ResNet experiments, we utilize imbalanced and/or long-tail versions of CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), SVHN (Netzer et al., 2011) and PathMNIST (Yang et al., 2021) datasets.
+For the LabelBench experiments, we utilize the FMoW (Christie et al., 2018), iWildcam (Beery et al., 2021), iNaturalist (Van Horn et al., 2018) and ImageNet-LT (Deng et al., 2009) datasets.
 We refer the readers to Appendix D for more details on our experiment setups.
-(a) Imbalanced CIFAR-10, three classes
-(b) Imbalanced CIFAR-100, ten classes
-(c) Imbalanced SVHN, two classes
-(d) PathMNIST, two classes
-(e) FMoW, Balanced Pool Accuracy
-(f) Imbalanced SVHN, two classes, # of minority labeled examples
-Figure 4: Performance comparison of DIRECT against other baselines algorithms in the noiseless but imbalanced setting.
-(a)-(d) are balanced accuracy on ResNet-18 experiments while (e) shows experiment under the LabelBench framework.
-$B_{\text{parallel}}$ indicates the number of parallel annotations as mentioned in Section 4.2.
-$B_{\text{parallel}}=1$ is equivalent with sychornous labeling.
+(a) CIFAR-10LT, 10 classes, 15% label noise
+(b) Imbalanced SVHN, two classes, 10% label noise
+(c) iWildcam Balanced Pool Accuracy, $10\%$ label noise
+Figure 5: Performance of DIRECT against baseline algorithms under label noise.
+(a)-(b) are balanced accuracy on ResNet-18 experiments while (c) shows results under the LabelBench framework.
 Results are averaged over four trials and the shaded areas represent standard errors around the mean.
+(a) Imbalanced CIFAR-100, two classes, no label noise
+(b) Imbalanced CIFAR-100, two classes, 10% label noise
+(c) Imbalanced CIFAR-100, two classes, 15% label noise
+Figure 6: Performance of DIRECT against baseline algorithms under different levels of label noise. Results are averaged over four trials and the shaded areas represent standard errors around the mean.
 5.1 Experiments under Imbalance, without Label Noise
@@
 The rest of the algorithms primarily optimizes for different notions of informativeness such as diversity and uncertainty.
+Figure 7: Performance of different active learning algorithm across different classes for the ImageNet-LT dataset.
 For the LabelBench experiments, due to the large dataset and model embedding sizes, we choose a subset of the algorithms that are computationally efficient and among top performers in the ResNet-18 results, including BADGE, Margin Sampling, CORESET and GALAXY.
@@
 Furthermore, in these experiments we also see using $B_{\text{parallel}}=5$ only affects algorithm performances minimally for DIRECT.
-(a) Imbalanced CIFAR-10, two classes, 10% label noise
-(b) Imbalanced SVHN, two classes, 10% label noise
-(c) iWildcam Balanced Pool Accuracy, $10\%$ label noise
-Figure 5: Performance of DIRECT against baseline algorithms under $10\%$ label noise.
-(a)-(b) are balanced accuracy on ResNet-18 experiments while (c) shows results under the LabelBench framework.
-Results are averaged over four trials and the shaded areas represent standard errors around the mean.
-(a) Imbalanced CIFAR-100, two classes, no label noise
-(b) Imbalanced CIFAR-100, two classes, 10% label noise
-(c) Imbalanced CIFAR-100, two classes, 15% label noise
-Figure 6: Performance of DIRECT against baseline algorithms under different levels of label noise. Results are averaged over four trials and the shaded areas represent standard errors around the mean.
 5.2 Experiments under Imbalance and Label Noise
@@
 For all of our experiments, we introduce a fixed percentage of label noise, where the given fraction of the examples’ labels are corrupted to a different class uniformly at random.
-For most of our experiements with $10\%$ label noise shown in Figure 5, we observe again that DIRECT consistently improves over all baselines including GALAXY.
+For most of our experiments with $10\%$ label noise shown in Figure 5, we observe again that DIRECT consistently improves over all baselines including GALAXY.
 The results are consistent on ResNet-18 and LabelBench setups, and with different $B_{\text{parallel}}$ values, showing DIRECT’s robustness under label noise.
@@
 In contrast, DIRECT significantly outperforms random sampling, saving more than $60\%$ of the annotation cost.
+5.3 Qualitative Analysis
+The ImageNet-LT dataset is constructed with class frequencies that gradually decrease according to class index.
+In Figure 7, we organize classes into bins ordered from most frequent to least frequent.
+The results show that DIRECT significantly outperforms baseline algorithms on less frequent classes (indices 301-1000), which accounts for DIRECT’s superior overall performance despite modest accuracy reductions on more frequent classes.
+This finding aligns with our balancedness analysis, where DIRECT demonstrates improved labeling of samples from rare classes.
 6 Conclusion and Future Work
@@
 We proposed an algorithm DIRECT that significantly and consistently outperforms existing literature.
-In this work, we also addressed the batch sampling problem of current SOTA algorithm, GALAXY (Zhang et al., 2022), by annotating multiple examples in parallel.
+In this work, we also addressed the batch sampling problem of Zhang et al.
+(2022), by annotating multiple examples in parallel.
 Studying asynchronous labeling could be a natural extension of our work.
@@
 In addition, one can further batch the labeling process across different classes to further accommodate an even larger number of parallel annotators.
-Lastly, we refer the readers to Appendix F for our impact statement.
+Impact Statement
+In the rapidly evolving landscape of machine learning, the efficacy of active learning in addressing data imbalance and label noise is a significant stride towards more robust and equitable AI systems.
+This research explores how active learning can effectively mitigate the challenges posed by imbalanced datasets and erroneous labels, prevalent in real-world scenarios.
+The positive impacts of this research are multifaceted.
+It enhances the accessibility and utility of machine learning in domains where data imbalance is a common challenge, such as healthcare, finance, and social media analytics.
+By improving class-balancedness in annotated sets, models trained on these datasets are less biased and more representative of real-world distributions, leading to fairer and more accurate outcomes.
+Additionally, this research contributes to reducing the time and cost associated with data annotation, which is particularly beneficial in fields where expert annotation is expensive or scarce.
+However, if not carefully implemented, active learning strategies could inadvertently introduce new biases or amplify existing ones, particularly in scenarios where the initial data is severely imbalanced or contains deeply ingrained biases.
+Furthermore, the advanced nature of these techniques may widen the gap between organizations with access to state-of-the-art technology and those without, potentially exacerbating existing inequalities in technology deployment.
 Acknowledgements
@@
 We further note that the time complexity factor of $K$ in DIRECT can be easily parallelized by conducting the $K$ sorting procedures on different CPU cores.
-In total, our experiments are conducted on NVIDIA 3090 ti GPUs. Each trial of the ResNet-18 experiment takes less than two hours while each trial of the LabelBench experiments takes roughly 12 hours.
+Below, we also provide a comprehensive list of computational complexity of different algorithms we consider in our implementation: As for data selection algorithms, let $K$ be the number of classes, $N$ be the pool size and $B_{\text{train}}$ be the batch size, $D$ be the penultimate layer embedding dimension and $T$ be the number of batches.
+Below, we detail the computation cost of data selection of each algorithm we consider.
+  • DIRECT: $O(T(KN\log N+B_{train}N))$.
+  • GALAXY: $O(T(KN\log N+B_{train}KN))$
+  • BADGE: $O(TB_{train}N(K+D))$
+  • Margin sampling/most likely positive/confidence sampling: $O(TKN)$
+  • Coreset: $O(T^{2}B_{train}ND)$
+  • SIMILAR: $O(TB_{train}ND)$
+  • Cluster margin: $O(N^{2}\log N+TN(K+\log N))$
+  • BASE: $O(TN(D+B_{train}))$
+Overall, our experiments are conducted on NVIDIA 3090 ti GPUs. Each trial of the ResNet-18 experiment takes less than two hours while each trial of the LabelBench experiments takes roughly 12 hours.
 Appendix D Experiment Setup
@@
 ResNet-18 Experiments.
-ResNet-18 with passive training has been the standard evaluation in existing deep active literature [Ash et al., 2019, Zhang et al., 2022].
-Our experiment setup utilizes the CIFAR-10, CIFAR-100 [Krizhevsky et al., 2009], SVHN [Netzer et al., 2011] and PathMNIST [Yang et al., 2021] image classification datasets.
+ResNet-18 with passive training has been the standard evaluation in existing deep active literature (Ash et al., 2019; Zhang et al., 2022).
+Our experiment setup utilizes the CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), SVHN (Netzer et al., 2011) and PathMNIST (Yang et al., 2021) image classification datasets.
 The original forms of these datasets are roughly balanced across 9, 10 or 100 classes.
@@
 Imbalance ratios are shown in Table 1.
-For neural network training, we utilize the standard passive training on labeled examples with cross entropy loss and Adam optimizer [Kingma and Ba, 2014].
-The ResNet-18 model [He et al., 2016] is pretrained on ImageNet [Deng et al., 2009] from the PyTorch library.
+In addition, we also utilize the standard CIFAR-10LT and CIFAR-100LT variants in our experiments for noisy label setting.
+For neural network training, we utilize the standard passive training on labeled examples with cross entropy loss and Adam optimizer (Kingma & Ba, 2014).
+The ResNet-18 model (He et al., 2016) is pretrained on ImageNet (Deng et al., 2009) from the PyTorch library.
 To address data imbalance, for all algorithms, we utilize a reweighted cross entropy loss by the inverse frequency of the number of labeled examples in each class.
-For experiments with label noise, we further add a $10\%$ label smoothing during training [Müller et al., 2019] for all algorithms.
+For experiments with label noise, we further add a $10\%$ label smoothing during training (Müller et al., 2019) for all algorithms.
 LabelBench Experiments.
 Proposed by Zhang et al.
-[2024a], LabelBench evaluates active learning performance in a more comprehensive framework.
-Here, we fine-tune the large pretrained model from CLIP’s ViT-B32 model [Radford et al., 2021].
-The framework also utilizes semi-supervised learning method FlexMatch [Zhang et al., 2021] to further leverage the unlabeled examples in the pool for training.
-We conduct experiments on the two imbalanced datasets in LabelBench, with FMoW [Christie et al., 2018] and iWildcam [Beery et al., 2021].
+(2024a), LabelBench evaluates active learning performance in a more comprehensive framework.
+Here, we fine-tune the large pretrained model from CLIP’s ViT-B32 model (Radford et al., 2021).
+The framework also utilizes semi-supervised learning method FlexMatch (Zhang et al., 2021) to further leverage the unlabeled examples in the pool for training.
+We conduct experiments on the two imbalanced datasets in LabelBench, with FMoW (Christie et al., 2018) and iWildcam (Beery et al., 2021).
 Similar to the ResNet-18 experiments, for all algorithms, we use a $10\%$ label smoothing in the loss function to improve training under label noise.
@@
 (b) iWildcam Balanced Pool Accuracy
+(c) ImageNet-LT Balanced Pool Accuracy
+(d) iNaturalist Balanced Pool Accuracy
 Figure 16: LabelBench results in the noiseless setting.
@@
 Figure 24: Imbalanced SVHN, three classes, 10% label noise.
+(a) CIFAR-10LT, 10 classes, 15% label noise.
+(b) CIFAR-100LT, 100 classes, 15% label noise.
 (a) FMoW Balanced Pool Accuracy
@@
 Figure 26: LabelBench results in the 10% label noise setting.
-Appendix F Impact Statement
-In the rapidly evolving landscape of machine learning, the efficacy of active learning in addressing data imbalance and label noise is a significant stride towards more robust and equitable AI systems.
-This research explores how active learning can effectively mitigate the challenges posed by imbalanced datasets and erroneous labels, prevalent in real-world scenarios.
-The positive impacts of this research are multifaceted.
-It enhances the accessibility and utility of machine learning in domains where data imbalance is a common challenge, such as healthcare, finance, and social media analytics.
-By improving class-balancedness in annotated sets, models trained on these datasets are less biased and more representative of real-world distributions, leading to fairer and more accurate outcomes.
-Additionally, this research contributes to reducing the time and cost associated with data annotation, which is particularly beneficial in fields where expert annotation is expensive or scarce.
-However, if not carefully implemented, active learning strategies could inadvertently introduce new biases or amplify existing ones, particularly in scenarios where the initial data is severely imbalanced or contains deeply ingrained biases.
-Furthermore, the advanced nature of these techniques may widen the gap between organizations with access to state-of-the-art technology and those without, potentially exacerbating existing inequalities in technology deployment.